### v3

# Whitney Approximation: domains and bounds

Abstract.
We investigate properties of holomorphic extensions in the one-variable case of Whitney’s Approximation Theorem on intervals. Improving a result of Gauthier-Kienzle, we construct tangentially approximating functions which extend holomorphically to domains of optimal size. For approximands on unbounded closed intervals, we also bound the growth of holomorphic extensions, in the spirit of Arakelyan, Bernstein, Keldych, and Kober.
1991 Mathematics Subject Classification:
30E10, 41A30
Introduction
The Weierstrass Approximation Theorem regarding the approximation of a continuous function $f$ on a compact interval by polynomials was generalized to the complex domain by Mergelyan [18] in 1951, who replaced the compact interval by a compact subset $K$ of the complex plane having connected complement with the additional (necessary) condition on $f$ that it be not only continuous on $K$ but also holomorphic on the interior of $K$. (See [12, Chapter III, §2] or [23, Chapter 20] for a proof.) This powerful theorem sparked the development of the subject of complex approximation theory in the second half of the 20th century, and has had numerous applications in many branches of complex analysis. Carleman [8] had previously (1927) found another analogue of the Weierstrass theorem, in which he replaced the compact interval by the real line and polynomials by entire functions. (Actually, Carleman’s theorem can be deduced without much effort from Mergelyan’s, cf. [21, proof of Theorem 8].) Staying with the approximation of functions on the real line, the present paper is devoted to a generalization of Carleman’s theorem due to Whitney (1934).
To explain Whitney’s result, we fix some conventions which will remain in force throughout the rest of this paper: we let $a$, $b$, $s$, $t$ range over $\mathbb{R}$}, $m$, $n$ over ${\mathbb{N}}=\{0,1,2,\dots\}$, $r$ over $\mathbb{N}$$\cup\{\infty\}$}, and $z$ over $\mathbb{C}$}. Let $I=(\alpha,\beta)$, where $\infty\leqslant\alpha<\beta\leqslant\infty$, be an open interval. Each $\mathcal{C}^{r}$-function $I\to{\mathbb{R}}$} and its derivatives can be arbitrarily well approximated by an analytic function and its derivatives; more precisely:
Theorem (Whitney).
Let $f\in\mathcal{C}^{r}(I)$ and $\varepsilon,\rho\in\mathcal{C}(I)$ be such that $\varepsilon>0$ and $r\geqslant\rho\geqslant 0$ on $I$. Then there is a $g\in\mathcal{C}^{\omega}(I)$ such that
(1) $\lvert(f-g)^{(n)}(t)\rvert<\varepsilon(t)\qquad\text{ for all~$t\in I$ and~$n\leqslant\rho(t)$.}$
This theorem is implicit in the proof of a (more general) lemma in Whitney’s seminal 1934 paper on extending multivariate differentiable functions [27]. In [3, Appendix A] we included a self-contained proof of this fact, which makes it apparent that there is an open subset $U=U_{I}$ of $\mathbb{C}$} which contains $I$, independent of $f$, such that the function $g$ in the conclusion of the theorem can be taken as the restriction of a holomorphic function $U\to{\mathbb{C}}$}. If $I$ is bounded, then the open set $U$ obtained through the proof of Whitney’s theorem as in loc. cit. is also bounded. (See [3, Corollary A.4].) However, if $I={\mathbb{R}}$, then this procedure yields $U={\mathbb{C}}$, so $g$ can be taken as the restriction of an entire function. Independently of Whitney, this fact had also been discovered by Hoischen [14]; for $r=0$, this is Carleman’s 1927 theorem [8] mentioned above, and for $r=1$ this was shown by Kaplan [15] in 1955. (See [21] for the history of such approximation theorems, going back to Weierstrass.) By a theorem of Gauthier and Kienzle [13, Theorem 1.3], in the case $r<\infty$} one can always take $U=I\cup({\mathbb{C}}\setminus{\mathbb{R}})$. Our first main result strengthens this fact by showing that it is possible to choose $U={\mathbb{C}}\setminus\{\alpha,\beta\}$ (even when $r=\infty$):
[…]
1
Theorem 1.
For all $f$, $\varepsilon$, $\rho$ as in Whitney’s Theorem, there is a $g\in\mathcal{C}^{\omega}(I)$ satisfying (1) which extends to a holomorphic function ${\mathbb{C}}\setminus\{\alpha,\beta\}\to{\mathbb{C}}$.
The proof of this theorem, given in Section 2 below, involves a simple reduction to Whitney’s theorem in the case $I={\mathbb{R}}$, by employing compositions with suitably chosen rational functions and some estimates involving Faà di Bruno’s Formula. In this section we also generalize Theorem 1 to other kinds of intervals:
[…]
1
Corollary 1.
Let $I\subseteq{\mathbb{R}}$ be connected and $f\in\mathcal{C}^{r}(I)$ (i.e., $f$ extends to a function in $\mathcal{C}^{r}(J)$ for some open interval $J\supseteq I$), and let $\varepsilon,\rho\in\mathcal{C}(I)$ be such that $\varepsilon>0$ and $r\geqslant\rho\geqslant 0$ on $I$. Then there is a $g\in\mathcal{C}^{\omega}(I)$ satisfying (1) which extends to a holomorphic function ${\mathbb{C}}\setminus\operatorname{fr}(I)\to{\mathbb{C}}$, where $\operatorname{fr}(I)=\operatorname{cl}(I)\setminus I$ is the frontier of $I$.
The second topic of this note are versions of Whitney’s theorem which include explicit bounds on the growth of the holomorphic extension of $g$ to $U$. The following example, due to Arakelyan [2, pp. 14–15], indicates why in the case of Carleman’s Theorem ($I={\mathbb{R}}$} and $r=0$) one cannot expect to bound the growth of the entire extension of $g$ solely from information about that of $f$:
Example.
Let $\beta\in\mathcal{C}({\mathbb{R}})$ be even and strictly increasing on $\mathbb{R}$$^{\geqslant}$ such that $\beta(0)=0$ and $\beta(t)\to\infty$} as $t\to\infty$, and let $\varepsilon\in\mathcal{C}({\mathbb{R}})$ be bounded with $\varepsilon>0$. Then the continuous function $f\colon{\mathbb{R}}\to{\mathbb{R}}$, $f(t):=\varepsilon(t)\cos\beta(t)$ is also bounded. Call an entire function $g$ real if $g({\mathbb{R}})\subseteq{\mathbb{R}}$. Suppose now that $g$ is a real entire function such that $\lvert f(t)-g(t)\rvert<\varepsilon(t)$ for each $t$. Then for each $t$ we have
$1+\cos\beta(t)<\frac{g(t)}{\varepsilon(t)}<1+\cos\beta(t),$
so between any two reals $a$, $b$ with $\cos\beta(a)=1$, $\cos\beta(b)=-1$ there is a zero of $g$. This implies that each interval [t,t][-t,t] ($t\geqslant 0$) contains $\geqslant(2/\pi)\beta(t)-2$ zeros of $g$. Using Jensen’s Formula [6, 1.2.1] we obtain a $C\in{\mathbb{R}}$ such that
$\log\,\lVert g\rVert_{t}\geqslant\frac{2}{\pi}\int_{1}^{t}\frac{\beta(s)}{s}\,ds-2\log t+C\qquad\text{for each $t\geqslant 1$.}$
Here and below $\lVert g\rVert_{t}:=\sup\big{\{}\lvert g(z)\rvert:\lvert z\rvert\leqslant t\big{\}}$ for $t\geqslant 0$.
However, Bernstein [5] and Kober [17] discovered that in this situation, for constant error function $\varepsilon$, the growth of the entire function $g$ can be limited, provided that $f$ is not only assumed to be bounded, but also uniformly continuous. (See [11, Theorem 2.1].) Keldych [16] proved a similar result, assuming that $f$ is $\mathcal{C}^{1}$ and the asymptotics of both $f$ and $f^{\prime}$ are known (see [12, p. 163]), and this was refined and extended to more general error functions by Arakelyan [1, 2]. Our second theorem provides a partial generalization of these facts to the case $r>0$, under some natural restrictions on $\varepsilon$$\rho$. To formulate it, we introduce some notation. Let $f\in\mathcal{C}^{m}({\mathbb{R}})$. For $t\geqslant 0$ and $\rho\in{\mathbb{R}}$ with $0\leqslant\rho\leqslant m$ put
$\lVert f\rVert_{t;\,\rho}:=\sup\big{\{}\lvert f^{(n)}(s)\rvert:\lvert s\rvert\leqslant t,\ n\leqslant\rho\big{\}}\in{\mathbb{R}}^{\geqslant},\qquad\lVert f\rVert_{t}:=\lVert f\rVert_{t;\,0}.$
Note that $\lVert f\rVert_{t_{1};\,\rho_{1}}\leqslant\lVert f\rVert_{t_{2};\,\rho_{2}}$ if $0\leqslant t_{1}\leqslant t_{2}$ and $0\leqslant\rho_{1}\leqslant\rho_{2}\leqslant m$. By convention we set $\infty+1:=\infty$, and for $\phi\colon{\mathbb{R}}^{\geqslant a}\to{\mathbb{R}}$ we let $\Delta\phi\colon{\mathbb{R}}^{\geqslant a}\to{\mathbb{R}}$ denote the difference function $t\mapsto\phi(t)-\phi(t+1)$. (If $\phi$ is strictly decreasing, then $\Delta\phi(t)>0$ for $t\geqslant a$, and if $\phi$ is convex, then $\Delta\phi$ is decreasing: see Section 1.) As usual $\log^{+}t:=\max\{0,\log t\}$, where $\log t:=$-$\infty$ for $t\leqslant 0$.
[…]
2
Theorem 2.
For each $f\in\mathcal{C}^{r+1}({\mathbb{R}})$} and $\varepsilon,\rho\in\mathcal{C}({\mathbb{R}}^{\geqslant})$ where $\varepsilon>0$ is strictly decreasing and convex with $\varepsilon(t)\to 0$} as $t\to\infty$, and $\rho$ with $r\geqslant\rho\geqslant 0$ is increasing, there are some $C,D\in{\mathbb{R}}^{>}$ and a real entire function $g$ such that
$\lvert(f-g)^{(n)}(t)\rvert<\varepsilon(\lvert t\rvert)\quad\text{ for $n\leqslant\rho(\lvert t\rvert)$,}$
and
$\lVert g\rVert_{t}\leqslant\exp\big{(}C\cdot s^{2}\cdot\big{(}1+\log^{+}\lVert f\rVert_{s+3\sqrt{2}}+\lambda(s+3\sqrt{2})\big{)}\big{)}\quad\text{for $t\geqslant 0$, $s=\sqrt{2}t+1$,}$
where
$\lambda(s):=\big{(}D(\rho(s)+1)\big{)}^{D(\rho(s)+1)s}\cdot\left(\frac{\lVert f\rVert_{s;\,\rho(s)+1}}{\Delta\varepsilon(s)}\right)^{3}.$
Here $C$ can be chosen independently of $\rho$, and $D$ independently of $f$, $\varepsilon$, $\rho$, $r$.
The proof of this theorem is given in Section 5, and is obtained by making the already quite constructive argument of Whitney completely explicit. This requires some bounds for derivatives of bump functions and for the approximation of functions of bounded support by Weierstrass transforms, computed in Sections 3 and 4 after some preliminaries in Section 1. In the rest of this introduction we collect some applications of the previous theorem, and also state a variant for functions $f\in\mathcal{C}^{r+1}({\mathbb{R}}^{\geqslant})$.
First, for $r<\infty$, by Theorem 2 applied to the constant function $\rho(t):=r$}:
[…]
2
Corollary 2.
[…]
$\lVert g\rVert_{t}\leqslant\exp\left(C\cdot s^{2}\cdot\left(1+\log^{+}\lVert f\rVert_{s}+D^{s}\left(\frac{\lVert f\rVert_{s;\,r+1}}{\Delta\varepsilon(s)}\right)^{3}\right)\right)$
for $t\geqslant 0$ and $s=\sqrt{2}(t+4)$. Here $D$ can be chosen to only depend on $r$.
For $r=0$ and certain $\varepsilon$, Arakelyan [2] has a qualitatively better bound:
Remark.
[…]
$\lVert g\rVert_{t}\leqslant\exp\left(C\cdot s\cdot\left(1+\log^{+}\left(\frac{\lVert f\rVert_{s}}{\varepsilon(t)}\right)+\frac{\lVert f^{\prime}\rVert_{s}}{\varepsilon(t)}\right)\right)\quad\text{if $t\geqslant 0$, $s=\sqrt{2}(t+4)$.}$
To facilitate the comparison with the bound in Corollary 2, note that if in addition $\varepsilon$ is assumed to be $\mathcal{C}^{2}$, strictly decreasing, and convex, then we have $1/\Delta\varepsilon(s)=O\big{(}{(s+1)/\varepsilon(s+1)}\big{)}$ as $s\to\infty$. (Lemma 1.2.)
Before stating the next corollary, we define some quantities measuring the growth of an entire function $g$ introduced in [24]. First recall that except in the case where $g$ is a constant, the function $t\mapsto\lVert g\rVert_{t}\colon{\mathbb{R}}^{\geqslant}\to{\mathbb{R}}^{\geqslant}$ is strictly increasing with $\lVert g\rVert_{t}\to\infty$ as $t\to\infty$. If $g$ is given by a polynomial of degree $n$ then $\lVert g\rVert_{t}=O(t^{n})$ as $t\to\infty$, and if $\lambda\in{\mathbb{R}}^{\geqslant}$ is such that $\lVert g\rVert_{t}=O(t^{\lambda})$ as $t\to\infty$, then $g$ is given by a polynomial of degree at most $\lfloor\lambda\rfloor$. Suppose now that $g$ is non-constant. With $\log_{m}$ denoting the $m$-fold iterated logarithm, define
$\lambda_{m}(g):=\limsup_{t\to\infty}\frac{\log_{m}\lVert g\rVert_{t}}{\log t}\in[0,\infty].$
If $\lambda_{m}(g)<\infty$ for some $m$, then $g$ is said to have finite index, and in this case the smallest such $m$ is called the index of $g$. Note that if $g$ has finite index $m$, then $m\geqslant 1$}. By convention, constant functions $\mathbb{C}$$\to{\mathbb{C}}$ have index 0. The entire functions of index $1$ are exactly the non-constant polynomial functions. The entire functions of index $\leqslant 2$ are also known as the entire functions of finite order, and  $\lambda_{2}(g)$ is called the order of $g$. (See [6, Chapter 2].) By [24, Theorem 1], if $g$ has index $m\geqslant 2$, then $\lambda=\lambda_{m}(g)$ may be computed from the Taylor coefficients $g_{n}:=g^{(n)}(0)/n!$ of $g$ at 0 as follows:
$\lambda=\limsup_{n\to\infty}\frac{n\log_{m-1}n}{-\log\lvert g_{n}\rvert}$.
For $f\in\mathcal{C}({\mathbb{R}})$ put $\lVert f\rVert:=\sup_{t\geqslant 0}\lVert f\rVert_{t}\in[0,\infty]$. From Corollary 2 we obtain:
[…]
3
Corollary 3.
Suppose $r<\infty$, and let $f\in\mathcal{C}^{r+1}({\mathbb{R}})$ be such that $\lVert f^{(n)}\rVert<\infty$ for $n\leqslant r+1$, and $\varepsilon\in{\mathbb{R}}^{>}$. Then there is a real entire function $g$ of index $\leqslant 3$ such that $\lvert({f-g})^{(n)}(t)\rvert<\varepsilon$ for $n\leqslant r$.
Proof.
Consider $\varepsilon_{0}\in\mathcal{C}({\mathbb{R}}^{\geqslant})$, $\varepsilon_{0}(t):=\varepsilon/(t+1)$; then $\Delta\varepsilon_{0}(t)=\frac{\varepsilon}{(t+1)(t+2)}$ for $t\geqslant 0$}. Take $C$, $D$, $g$ as in Corollary 2 applied to $\varepsilon_{0}$ in place of $\varepsilon$. Then with $M:=\max\!\big{\{}1,\lVert f\rVert_{0},\dots,\lVert f\rVert_{r+1}\big{\}}$, for $t\geqslant 0$ and $s=\sqrt{2}(t+4)$ we have
$\lVert g\rVert_{t}\leqslant\exp\left(C\cdot s^{2}\cdot\left(1+\log M+D^{s}\left(\frac{M(s+1)(s+2)}{\varepsilon}\right)^{3}\right)\right).$
For suitable $E\in{\mathbb{R}}^{\geqslant 1}$ (only depending on $C$, $D$, $\varepsilon$, $M$) we thus have $\lVert g\rVert_{t}\leqslant\exp(E^{s})$ for all sufficiently large $t\geqslant 0$, and this yields $\lambda_{3}(g)\leqslant 1$. ∎
Remarks.
Let $f\in\mathcal{C}^{1}({\mathbb{R}})$ and $M\in{\mathbb{R}}^{\geqslant 1}$ be such that $\lVert f\rVert,\lVert f^{\prime}\rVert\leqslant M$, and $\varepsilon\in{\mathbb{R}}^{>}$. Then by the remark following Corollary 2 (applied to the constant function $\varepsilon$), there is a $C\in{\mathbb{R}}^{>}$ and a real entire function $g$ such that $\lvert(f-g)(t)\rvert<\varepsilon$ for all $t$ and
$\lVert g\rVert_{t}\leqslant\exp\left(C\cdot s\cdot\left(1+\log^{+}\left(\frac{M}{\varepsilon}\right)+\frac{M}{\varepsilon}\right)\right)\quad\text{if $t\geqslant 0$, $s=\sqrt{2}(t+4)$,}$
thus $\lambda_{2}(g)\leqslant 1$. Therefore in Corollary 3, in the case $r=0$ one can replace “index $\leqslant 3$” by “order $\leqslant 1$”. (This is originally due to Bernstein [5] and Kober [17].) It would be interesting to know whether in the case $r>0$, one can improve “index $\leqslant 3$” to “finite order” in the previous corollary.
The case $r=\infty$ of Theorem 2 leads to an application to rapidly decreasing functions: Let $f\in\mathcal{C}^{\infty}({\mathbb{R}})$, and put $\lVert f\rVert_{m,n}:=\sup\big{\{}\lvert t^{m}f^{(n)}(t)\rvert:t\in{\mathbb{R}}\big{\}}\in[0,\infty]$. Then $f$ is said to be rapidly decreasing if $\lVert f\rVert_{m,n}<\infty$ for all $m$, $n$; equivalently, if for all $m$, $n$ we have $t^{m}f^{(n)}(t)\to 0$ as $\lvert t\rvert\to\infty$. Examples: if $f$ has bounded support, then it is rapidly decreasing; and for each $\lambda\in{\mathbb{R}}^{>}$ and $m$, the function $t\mapsto t^{m}\operatorname{e}^{-\lambda t^{2}}\colon{\mathbb{R}}\to{\mathbb{R}}$ is rapidly decreasing. The rapidly decreasing functions form a subalgebra $\mathcal{S}({\mathbb{R}})$ of the $\mathbb{R}$}-algebra $\mathcal{C}^{\infty}({\mathbb{R}})$. The collection of sets
$\big{\{}g\in\mathcal{S}({\mathbb{R}}):\text{$\lVert f-g\rVert_{m,n}<\varepsilon$ for $m,n\leqslant N$}\big{\}}\quad\text{where $f\in\mathcal{S}({\mathbb{R}})$, $\varepsilon\in{\mathbb{R}}^{>}$, and $N\in{\mathbb{N}}$,}$
is the basis of a topology on $\mathcal{S}({\mathbb{R}})$, which makes the $\mathbb{R}$}-vector space $\mathcal{S}({\mathbb{R}})$ into a topological $\mathbb{R}$}-vector space (in fact, a Fréchet space), known as the Schwartz space on $\mathbb{R}$}, which plays an important role in the theory of distributions (cf. [25]). It is well-known that the set of rapidly decreasing functions which extend to entire functions is dense in $\mathcal{S}({\mathbb{R}})$; see [25, Theorem 15.5, p. 160]. From Theorem 2 we obtain an effective version of this fact:
[…]
4
Corollary 4.
Let $f\in\mathcal{S}({\mathbb{R}})$, $\varepsilon\in{\mathbb{R}}^{>}$, and $N\in{\mathbb{N}}$. Then there is a real entire function $g$ such that $g|_{{\mathbb{R}}}\in\mathcal{S}({\mathbb{R}})$ and $\lVert f-g\rVert_{m,n}\leqslant\varepsilon$ for each $m,n\leqslant N$. Moreover, we can choose $g$ so that for some $C,D\in{\mathbb{R}}^{\geqslant 1}$ we have
$\lVert\widehat{g}\rVert_{t}\leqslant\exp\big{(}C\cdot s^{2}\cdot\big{(}1+\log^{+}\lVert f\rVert_{s+3\sqrt{2}}+\lambda(s+3\sqrt{2})\big{)}\big{)}\quad\text{for $t\geqslant 0$, $s=\sqrt{2}t+1$,}$
where
$\lambda(s):=\big{(}D\,(r+1)\big{)}^{D\,(r+1)\,s}\cdot\left(\frac{(N/\!\operatorname{e})^{N}\lVert f\rVert_{0,\lceil r+1\rceil}}{\varepsilon}\right)^{3}\qquad\text{where $r=N+s$.}$
Here $C$ can be chosen independently of $N$, and $D$ independently of $f$, $\varepsilon$, $N$.
Proof.
Put $\delta:=\varepsilon/(N/\!\operatorname{e})^{N}\in{\mathbb{R}}^{>}$ and consider $\varepsilon_{0},\rho_{0}\in\mathcal{C}({\mathbb{R}}^{\geqslant})$ given by $\varepsilon_{0}(t):=\delta\operatorname{e}^{-t}$ and $\rho_{0}(t):=N+t$ for $t\geqslant 0$. Take $g$ as in Theorem 2 applied to $\varepsilon_{0}$, $\rho_{0}$ in place of $\varepsilon$$\rho$. Then for each $m$, $n$ we have
$\lvert t^{m}g^{(n)}(t)\rvert\leqslant\lvert t^{m}f^{(n)}\rvert+\delta\lvert t\rvert^{m}\operatorname{e}^{-\lvert t\rvert}\quad\text{if $\lvert t\rvert\geqslant n-N$,}$
thus
$\lVert g\rVert_{m,n}\leqslant\lVert x^{m}g^{(n)}\rVert_{\max\{0,n-N\}}+\lVert f\rVert_{m,n}+\delta(m/\!\operatorname{e})^{m}<\infty.$
Hence $g|_{{\mathbb{R}}}\in\mathcal{S}({\mathbb{R}})$; moreover, if $m,n\leqslant N$ then
$\lvert t^{m}(f-g)^{(n)}\rvert\leqslant\delta\lvert t\rvert^{m}\operatorname{e}^{-\lvert t\rvert}\leqslant\delta\lvert t\rvert^{N}\operatorname{e}^{-\lvert t\rvert}\leqslant\varepsilon\quad\text{for each $t$,}$
hence $\lVert f-g\rVert_{m,n}\leqslant\varepsilon$. The rest now follows from Theorem 2. Note that $\Delta\varepsilon_{0}(t)=\delta\operatorname{e}^{-t}(1-\operatorname{e}^{-1})\geqslant\varepsilon_{0}(t)/2$ and so $1/\Delta\varepsilon_{0}(t)\leqslant(2/\varepsilon)(N/\!\operatorname{e})^{N}\operatorname{e}^{t}$, for each $t\geqslant 0$. ∎
Finally, in Section 5 we also establish a version of Theorem 2 for $f\in\mathcal{C}^{r+1}({\mathbb{R}}^{\geqslant})$. For such $f$, given $\varepsilon,\rho\in\mathcal{C}({\mathbb{R}}^{\geqslant})$ with $\varepsilon>0$ and $r\geqslant\rho\geqslant 0$, Corollary 1 yields a real entire function $g$ such that $\lvert(f-g)^{(n)}(t)\rvert<\varepsilon(t)$ for $n\leqslant\rho(t)$, $t\geqslant 0$. Here we will confine ourselves to constructing a holomorphic approximation of $f$ with domain $U:=\big{\{}z\in{\mathbb{C}}:\lvert\operatorname{Im}z\rvert<\operatorname{Re}z\big{\}}$ while at the same time controlling its growth, as in Theorem 2, on a proper subset of $U$. Notation: for $f\in\mathcal{C}^{m}({\mathbb{R}}^{\geqslant})$, $t\geqslant 0$, and $\rho\in[0,m]$ put
$\lVert f\rVert_{t;\,\rho}:=\sup\big{\{}\lvert f^{(n)}(s)\rvert:0\leqslant s\leqslant t,\ n\leqslant\rho\big{\}}\in{\mathbb{R}}^{\geqslant},\qquad\lVert f\rVert_{t}:=\lVert f\rVert_{t;\,0}.$
[…]
3
Theorem 3.
Let $\alpha\in{\mathbb{R}}^{>}$, $f\in\mathcal{C}^{r+1}({\mathbb{R}}^{\geqslant})$, and $\varepsilon,\rho\in\mathcal{C}({\mathbb{R}}^{>})$ where $\varepsilon>0$ is strictly decreasing and convex with $\varepsilon(t)\to 0$} as $t\to\infty$, and $\rho$ with $r\geqslant\rho\geqslant 0$ is increasing. Then there are some $C,D\in{\mathbb{R}}^{>}$ and a holomorphic $g\colon U\to{\mathbb{C}}$ such that $g({\mathbb{R}}^{>})\subseteq{\mathbb{R}}$,
$\lvert(f-g)^{(n)}(t)\rvert<\varepsilon(t)\quad\text{ for $n\leqslant\rho(t)$ and all $t>0$,}$
and for all $z\in{\mathbb{C}}$ with $\operatorname{Re}z>0$ and $(\operatorname{Im}z)^{2}\leqslant(\operatorname{Re}z)^{2}-\alpha$ we have
$\lvert g(z)\rvert\leqslant\exp\big{(}C\cdot s\cdot\big{(}1+\log^{+}\lVert f\rVert_{s}+\lambda(s)\big{)}\big{)}\quad\text{for $t\geqslant\lvert z\rvert$, $s=D(t^{2}+1)$,}$
where
$\lambda(s):=\big{(}D(\rho(s)+1)\big{)}^{D(\rho(s)+1)s^{2}}\cdot\left(\frac{\lVert f\rVert_{s;\,\rho(s)+1}}{\Delta\varepsilon(s)}\right)^{3}.$
Here $C$ can be chosen independently of $\rho$, and $D$ independently of $f$, $\varepsilon$, $\rho$, $r$.
Notations and conventions
Let $U\subseteq{\mathbb{R}}$} be open and $\emptyset\neq S\subseteq U$. For $f\in\mathcal{C}(U)$ we set
$\lVert f\rVert_{S}\ :=\ \sup\big{\{}\lvert f(s)\rvert:\,s\in S\big{\}}\in[0,\infty],$
so for $f,g\in\mathcal{C}(U)$ and $\lambda\in{\mathbb{R}}$ (and the convention $0\cdot\infty=\infty\cdot 0=0$) we have
$\lVert f+g\rVert_{S}\leqslant\lVert f\rVert_{S}+\lVert g\rVert_{S},\quad\lVert\lambda f\rVert_{S}=|\lambda|\cdot\lVert f\rVert_{S},\quad\text{and}\quad\lVert fg\rVert_{S}\leqslant\lVert f\rVert_{S}\lVert g\rVert_{S}.$
If $\emptyset\neq S^{\prime}\subseteq S$ then $\lVert f\rVert_{S^{\prime}}\leqslant\lVert f\rVert_{S}$. Next, let $f\in\mathcal{C}^{m}(U)$. We then put
$\lVert f\rVert_{S;\,m}\ :=\ \max\big{\{}\lVert f\rVert_{S},\dots,\lVert f^{(m)}\rVert_{S}\big{\}}\in[0,\infty].$
Then again for $f,g\in\mathcal{C}^{m}(U)$ and $\lambda\in{\mathbb{R}}$ we have
$\lVert f+g\rVert_{S;m}\leqslant\lVert f\rVert_{S;m}+\lVert g\rVert_{S;m},\quad\lVert\lambda f\rVert_{S;m}=|\lambda|\cdot\lVert f\rVert_{S;m},$
and
(0.1) $\lVert fg\rVert_{S;\,m}\ \leqslant\ 2^{m}\lVert f\rVert_{S;\,m}\lVert g\rVert_{S;\,m}.$
Also note that for $f\in\mathcal{C}^{m}({\mathbb{R}})$, $t\geqslant 0$, and $\rho\in[0,m]$ we have $\lVert f\rVert_{t;\,\rho}=\lVert f\rVert_{[-t,t];\,n}$ where $n=\lfloor r\rfloor$.
Let $f\in\mathcal{C}^{m}(U)$. For $U={\mathbb{R}}$ we set $\lVert f\rVert_{m}:=\lVert f\rVert_{{\mathbb{R}};\,m}$ and $\lVert f\rVert:=\lVert f\rVert_{0}$. For $k\leqslant m$ and $\emptyset\neq S^{\prime}\subseteq S\subseteq U$ we have $\lVert f\rVert_{S^{\prime};\,k}\leqslant\lVert f\rVert_{S;\,m}$. Moreover, $\lVert f\rVert_{S;\,m}$ does not change if $S$ is replaced by its closure in $U$.
Acknowledgements
The author would like to thank Paul Gauthier and Herwig Hauser for various suggestions to improve this paper.
1. Preliminaries
In this section we collect various auxiliary results used later in the paper: some remarks on difference functions of convex functions, estimates concerning factorials, and an easy bound for a sum of square roots.
The difference function of a convex function
In the next lemma we let $I$ be an interval (of any kind) in $\mathbb{R}$}, and we let $\phi\colon I\to{\mathbb{R}}$ be a convex function, that is,
$\phi\big{(}(1-\lambda)s+\lambda t\big{)}\leqslant(1-\lambda)\phi(s)+\lambda\phi(t)\quad\text{ for each~$\lambda\in[0,1]$ and $s,t\in I$.}$
We consider the function
$\Delta\colon D:=\big{\{}(s,t)\in I\times I:s\neq t\big{\}}\to{\mathbb{R}},\qquad\Delta(s,t):=\frac{\phi(s)-\phi(t)}{s-t}.$
Note that $D$ and $\Delta$ are symmetric. The following fact is well-known:
Lemma $1.$
Let $s<t<u$ be in $I$; then $\Delta(s,t)\leqslant\Delta(s,u)\leqslant\Delta(t,u)$.
Proof.
We have $t=\lambda s+\mu u$ where $\lambda:=(u-t)/(u-s)$ and $\mu:=1-\lambda=(t-s)/(u-s)$. Then $\phi(t)\leqslant\lambda\phi(s)+\mu\phi(u)$ by convexity of $\phi$. Subtracting $\phi(s)$ from both sides of this inequality yields
$\phi(t)-\phi(s)\leqslant\mu\big{(}\phi(s)-\phi(u)\big{)}=(t-s)\Delta(s,u),$
and so we get $\Delta(s,t)\leqslant\Delta(s,u)$, whereas subtracting $\phi(t)+\lambda\big{(}\phi(s)-\phi(u)\big{)}$ from both sides of this inequality yields
$(u-t)\Delta(s,u)=\lambda\big{(}\phi(u)-\phi(s)\big{)}\leqslant\phi(u)-\phi(t)$
and so $\Delta(s,u)\leqslant\Delta(t,u)$. ∎
Suppose now that $I$ is not bounded from above. Then by Lemma 1.1, the function
$t\mapsto\Delta\phi(t):=-\Delta(t,t+1)=\phi(t)-\phi(t+1)\colon I\to{\mathbb{R}}$
is decreasing. Also note that if $\phi\in\mathcal{C}^{2}(I)$, then $\phi^{\prime\prime}\geqslant 0$ by convexity of $\phi$, so $\Delta\phi(t)\geqslant-\phi^{\prime}(t+1)$ for $t\in I$, by the Mean Value Theorem. This yields:
Lemma 1.2.
Suppose $\phi\in\mathcal{C}^{2}(I)$, $\phi>0$, and $C\in{\mathbb{R}}^{>}$ are such that $\phi^{\prime}(t)/\phi(t)\leqslant-C/t$ for each $t\in I$ with $t>0$. Then $\Delta\phi(t)\geqslant C\phi(t+1)/(t+1)$ for $t\in I$, $t>0$.
Cheap approximations to the factorial
The following estimates for $n!$ are not as tight as Stirling’s approximation $n!\sim\sqrt{2\pi n}(n/{\operatorname{e}})^{n}$ (as $n\to\infty$), but good enough for our purposes:
Lemma 1.3.
Suppose $n\geqslant 1$. Then
$\operatorname{e}\left(\frac{n}{\operatorname{e}}\right)^{n}\leqslant n!\leqslant\frac{\operatorname{e}^{2}}{4}\left(\frac{n+1}{\operatorname{e}}\right)^{n+1}.$
Proof.
For $t>1$, from the inequalities $t-1\leqslant\lfloor t\rfloor\leqslant t$ we obtain $\log(t-1)\leqslant\log\lfloor t\rfloor\leqslant\log t$. Hence by integration
$\int_{2}^{n+1}\log(t-1)\,dt\leqslant\int_{2}^{n+1}\log\lfloor t\rfloor\,dt=\sum_{k=2}^{n}\log k\leqslant\int_{2}^{n+1}\log t\,dt$
and thus
$n\big{(}\log(n)-1\big{)}+1\leqslant\log(n!)\leqslant(n+1)\big{(}\log(n+1)-1\big{)}-\big{(}2\log(2)-2\big{)},$
and this yields the claimed inequality. ∎
Corollary 1.4.
Let $\rho\in{\mathbb{R}}$ with $\rho\geqslant\operatorname{e}t>0$. Then
$t^{\rho}\frac{4}{\operatorname{e}^{2}}\left(\frac{\operatorname{e}}{\rho+2}\right)^{\rho+2}\leqslant\frac{t^{n}}{n!}\qquad\text{for $n\leqslant\rho$.}$
Proof.
Put $m:=\lceil\rho\rceil$. Then $0\leqslant m-1<\rho\leqslant m$ and thus
(1.1) $\frac{4}{\operatorname{e}^{2}}\left(\frac{\operatorname{e}}{\rho+2}\right)^{\rho+2}\leqslant\frac{4}{\operatorname{e}^{2}}\left(\frac{\operatorname{e}}{m+1}\right)^{m+1}\leqslant\frac{1}{m!},$
where the second inequality holds by Lemma 1.3. Since $m\geqslant\operatorname{e}t$, by Lemma 1.3 again:
$\frac{t^{m}}{m!}\leqslant t^{m}\frac{1}{\operatorname{e}}\left(\frac{\operatorname{e}}{m}\right)^{m}\leqslant\left(\frac{\operatorname{e}t}{m}\right)^{m}\leqslant 1.$
Moreover, $t^{n}/n!\leqslant t^{n+1}/(n+1)!$ for $n\leqslant t-1$, and $t^{n}/n!\geqslant t^{n+1}/(n+1)!$ for $n\geqslant t-1$. Hence $\min\{t^{n}/n!:n\leqslant m\}=t^{m}/m!$, so in particular,
(1.2) $t^{m}/m!\leqslant t^{n}/n!\quad\text{ for $n\leqslant\rho$.}$
Now if $t\leqslant 1$ then for $n\leqslant\rho$ we have $t^{\rho}\leqslant t^{n}$ and thus by (1.1):
$t^{\rho}\frac{4}{\operatorname{e}^{2}}\left(\frac{\operatorname{e}}{\rho+2}\right)^{\rho+2}\leqslant\frac{t^{n}}{m!}\leqslant\frac{t^{n}}{n!}.$
If $t>1$, then $t^{\rho}\leqslant t^{m}$ and so
$t^{\rho}\frac{4}{\operatorname{e}^{2}}\left(\frac{\operatorname{e}}{\rho+2}\right)^{\rho+2}\leqslant\frac{t^{m}}{m!}\leqslant\frac{t^{n}}{n!},$
using both (1.1) and (1.2). ∎
Bounding a sum of square roots
For all $\alpha,\beta\in{\mathbb{R}}^{\geqslant}$ we have
$\sqrt{\alpha+\beta}\leqslant\sqrt{\alpha}+\sqrt{\beta}\leqslant\sqrt{2(\alpha+\beta)}.$
This observation easily yields the following fact, recorded here for later use:
Lemma 1.5.
If $a,b,c\in{\mathbb{R}}^{\geqslant}$, then $\sqrt{a+b}+\sqrt{c}\leqslant\sqrt{a}+\sqrt{2(b+c)}$.
2. Proof of Theorem 1
We begin with a general estimate coming out of a case of Faà di Bruno’s Formula [9, §3.4, Theorem B]. For this, let $f\in\mathcal{C}^{\infty}(I)$ where $I$ is an open interval in $\mathbb{R}$} and $g\in\mathcal{C}^{r}(J)$} where $J$ is an open interval with $f(I)\subseteq J$, and set $h:=g\circ f\in\mathcal{C}^{r}(I)$. Let $B_{mn}(y_{1},\dots,y_{n-m+1})\in{\mathbb{Q}}[y_{1},\dots,y_{n-m+1}]$ ($m\leqslant n$) be the Bell polynomials as defined in [4, 12.5]. Then for $n\leqslant r$ we have
(2.1) $h^{(n)}=\sum_{m=0}^{n}(g^{(m)}\circ f)\cdot B_{mn}(f^{\prime},f^{\prime\prime},\dots,f^{(n-m+1)}).$
[…]
Lemma 2.1.
Suppose $1\leqslant n\leqslant r$ and $t\in I$, and let $F,G\in{\mathbb{R}}$, $F\geqslant 1$ be such that
$\lvert f^{(k)}(t)\rvert\leqslant F\quad\text{ for $k=1,\dots,n$,}\qquad\lvert g^{(m)}(f(t))\rvert\leqslant G\quad\text{ for $m=1,\dots,n$.}$
Then $\lvert h^{(n)}(t)\rvert\leqslant G\cdot(nF)^{n}$.
Proof.
By (2.1) we have
$\lvert h^{(n)}(t)\rvert$ $\leqslant\sum_{m=1}^{n}G\cdot\lvert B_{mn}(f^{\prime}(t),f^{\prime\prime}(t),\dots,f^{(n-m+1)}(t))\rvert$
$\leqslant\sum_{m=1}^{n}G\cdot B_{mn}(1,1,\dots,1)\cdot F^{m}\leqslant G\cdot B_{n}\cdot F^{n}\leqslant G\cdot(nF)^{n}.\qed$
We now prove Theorem 1. Thus let $I$ be an open interval, $f\in\mathcal{C}^{r}(I)$, and $\varepsilon,\rho\in\mathcal{C}(I)$} where $\varepsilon>0$, $r\geqslant\rho\geqslant 0$. We first assume that $I$ is bounded, say $I=(a,b)$ where $a<b$, and we need to show the existence of a holomorphic function $g\colon{\mathbb{C}}\setminus\{a,b\}\to{\mathbb{C}}$ such that $g(I)\subseteq{\mathbb{R}}$ and $\lvert({f-g})^{(n)}(t)\rvert<\varepsilon(t)$ for all $t\in I$ and $n\leqslant\rho(t)$. It is straightforward to arrange that I=(1,1)I=(-1,1): Set $\alpha:=\frac{b-a}{2}$, $\beta:=\min\{\alpha,1\}$, and J:=(1,1)J:=(-1,1), and consider the holomorphic bijection $\phi\colon{\mathbb{C}}\to{\mathbb{C}}$, $\phi(z):=\alpha\cdot z+\textstyle\frac{1}{2}(a+b)$, with compositional inverse $\phi^{\operatorname{inv}}$. Put $f_{0}:=(f\circ\phi)|_{J}\colon J\to{\mathbb{R}}$ and let $\varepsilon_{0},\rho_{0}\in\mathcal{C}(J)$ be given by $\varepsilon_{0}(s):=\beta^{\rho(\phi(s))}\varepsilon(\phi(s))$ and $\rho_{0}(s):=\rho(\phi(s))$, respectively. Suppose we have a holomorphic function $g_{0}\colon{\mathbb{C}}\setminus\{\pm 1\}\to{\mathbb{C}}$ such that $g_{0}(s)\in{\mathbb{R}}$ and
$\lvert(f_{0}-g_{0})^{(n)}(s)\rvert<\varepsilon_{0}(s)\quad\text{for $s\in J$, $n\leqslant\rho_{0}(s)$.}$
Consider the holomorphic function $g:=g_{0}\circ\phi^{\operatorname{inv}}\colon{\mathbb{C}}\setminus\{a,b\}\to{\mathbb{C}}$. Then for $t\in I$ and $n\leqslant\rho(t)$, setting $s:=\phi^{\operatorname{inv}}(t)\in J$ we have $g(t)=g_{0}(s)\in{\mathbb{R}}$ and
$(f_{0}-g_{0})^{(n)}(s)=\big{(}(f-g)\circ\phi\big{)}^{(n)}(s)=(f-g)^{(n)}(t)\cdot\alpha^{n}$
and so
$\lvert(f-g)^{(n)}(t)\rvert<\beta^{\rho(t)}\alpha^{-n}\varepsilon(t)\leqslant\varepsilon(t).$
Hence replacing $f$, $\varepsilon$, $\rho$ by $f_{0}$, $\varepsilon_{0}$, $\rho_{0}$, respectively, we may arrange I=(1,1)I=(-1,1), which we assume from now on. The rational function
$\Phi\colon{\mathbb{C}}\setminus\{\pm 1\}\to{\mathbb{C}},\qquad\Phi(z):=\frac{2z}{1-z^{2}}=\frac{1}{1-z}-\frac{1}{z+1}$
restricts to an analytic bijection $\phi\colon I\to{\mathbb{R}}$. Let $\phi^{\operatorname{inv}}\colon{\mathbb{R}}\to I$} be its (analytic) compositional inverse. An easy induction on $n$ shows that for each $n$ and $t\in I$,
$\phi^{(n)}(t)=n!\left(\frac{1}{(1-t)^{n+1}}+(-1)^{n+1}\frac{1}{(t+1)^{n+1}}\right)$
and thus
(2.2) $\lvert\phi^{(n)}(t)\rvert\leqslant n!\frac{(t+1)^{n+1}+(1-t)^{n+1}}{(1-t^{2})^{n+1}}\leqslant\frac{n!\cdot 2^{n+1}}{(1-t^{2})^{n+1}}.$
Consider the function $\varepsilon_{*}\in\mathcal{C}(I)$ given by
$\varepsilon_{*}(t):=\varepsilon(t)\cdot\left(\frac{4}{(\rho(t)+1)\operatorname{e}^{2}}\cdot\left(\frac{\operatorname{e}(1-t^{2})}{2(\rho(t)+1)}\right)^{\rho(t)+1}\right)^{\rho(t)}\quad\text{for $t\in I$.}$
Whitney’s Theorem (in the case $I={\mathbb{R}}$) applied to $f_{*}:=f\circ\phi^{\operatorname{inv}}\in\mathcal{C}^{r}({\mathbb{R}})$ as well as ${\varepsilon_{*}\circ\phi^{\operatorname{inv}}},\rho\circ\phi^{\operatorname{inv}}\in\mathcal{C}({\mathbb{R}})$ in place of $f$, $\varepsilon$, $\rho$, respectively, yields an entire function $g_{*}\colon{\mathbb{C}}\to{\mathbb{C}}$ such that $g_{*}({\mathbb{R}})\subseteq{\mathbb{R}}$ and, with $h_{*}:=f_{*}-g_{*}|_{{\mathbb{R}}}$:
(2.3) $\lvert h_{*}^{(n)}(s)\rvert<\varepsilon_{*}(\phi^{\operatorname{inv}}(s))\qquad\text{ whenever $n\leqslant\rho(\phi^{\operatorname{inv}}(s))$.}$
Consider the holomorphic function $g:=g_{*}\circ\Phi\colon{\mathbb{C}}\setminus\{\pm 1\}\to{\mathbb{C}}$ and put $h:=f-g|_{I}$, so $h_{*}=h\circ\phi$. Then for $t\in I$ we have $\lvert h(t)\rvert<\varepsilon(t)$, and if $1\leqslant n\leqslant\rho(t)$, then
$\lvert h^{(n)}(t)\rvert$ $<\varepsilon_{*}(t)\cdot\left(\frac{n\cdot n!\cdot 2^{n+1}}{(1-t^{2})^{n+1}}\right)^{n}$
$\leqslant\varepsilon_{*}(t)\cdot\left(\frac{n\operatorname{e}^{2}}{4}\cdot\left(\frac{2(n+1)}{\operatorname{e}(1-t^{2})}\right)^{n+1}\right)^{n}$
$\leqslant\varepsilon_{*}(t)\cdot\left(\frac{\big{(}\rho(t)+1\big{)}\operatorname{e}^{2}}{4}\cdot\left(\frac{2(\rho(t)+1)}{\operatorname{e}(1-t^{2})}\right)^{\rho(t)+1}\right)^{\rho(t)}=\varepsilon(t),$
using (2.2) and (2.3) in combination with Lemma 2.1 for the first inequality and Lemma 1.3 for the second inequality.
This shows Theorem 1 in the case of a bounded interval. Now suppose $I\subseteq{\mathbb{R}}$ is unbounded, say $I=(a,\infty)$. Replacing $a$, $f$, $\varepsilon$, $\rho$ by 0, $t\mapsto f(t+a)$, $t\mapsto\varepsilon(t+a)$, and $t\mapsto\rho(t+a)$, respectively, we first arrange $a=0$, so $I={\mathbb{R}}^{>}$. Now consider the rational function
$\Psi\colon{\mathbb{C}}\setminus\{0\}\to{\mathbb{C}},\qquad\Psi(z):=z-\frac{1}{z},$
which restricts to an analytic bijection $\psi\colon{\mathbb{R}}^{>}\to{\mathbb{R}}$ with (analytic) compositional inverse $\psi^{\operatorname{inv}}\colon{\mathbb{R}}\to{\mathbb{R}}^{>}$. For $t>0$ we have
$\psi^{\prime}(t)=1+\frac{1}{t^{2}},\qquad\psi^{(n)}(t)=\frac{n!}{(-t)^{n+1}}\quad\text{if $n>1$.}$
For $t>0$, putting
$\rho_{+}(t):=\rho(t)+\operatorname{e}t,\qquad\beta(t):=1+t^{-(\rho_{+}(t)+1)}\frac{\operatorname{e}^{2}}{4}\left(\frac{\rho_{+}(t)+2}{\operatorname{e}}\right)^{\rho_{+}(t)+2}$
we obtain $\lvert\psi^{(n)}(t)\rvert\leqslant\beta(t)$ for $1\leqslant n\leqslant\rho_{+}(t)$, by Corollary 1.4. Consider
$\varepsilon_{*}\in\mathcal{C}({\mathbb{R}}^{>}),\qquad\varepsilon_{*}(t):=\varepsilon(t)\cdot\big{(}\rho(t)\beta(t)\big{)}^{-\rho(t)}\quad\text{for $t>0$.}$
As in the previous case, Whitney’s Theorem applied to $f_{*}:=f\circ\psi^{\operatorname{inv}}\in\mathcal{C}^{r}({\mathbb{R}})$ as well as ${\varepsilon_{*}\circ\psi^{\operatorname{inv}}},\rho\circ\psi^{\operatorname{inv}}\in\mathcal{C}({\mathbb{R}})$ in place of $f$, $\varepsilon$, $\rho$, respectively, yields an entire function $g_{*}\colon{\mathbb{C}}\to{\mathbb{C}}$ such that $g_{*}({\mathbb{R}})\subseteq{\mathbb{R}}$ and, with $h_{*}:=f_{*}-g_{*}|_{{\mathbb{R}}}$:
$\lvert h_{*}^{(n)}(s)\rvert<\varepsilon_{*}(\psi^{\operatorname{inv}}(s))\qquad\text{ whenever $n\leqslant\rho(\psi^{\operatorname{inv}}(s))$.}$
Let $g:=g_{*}\circ\Psi$, a holomorphic function ${\mathbb{C}}\setminus\{0\}\to{\mathbb{C}}$, and $h:=f-g|_{{\mathbb{R}}^{>}}\in\mathcal{C}^{r}({\mathbb{R}}^{>})$, so $h_{*}=h\circ\psi$. Then for $t>0$ we have $\lvert h(t)\rvert<\varepsilon(t)$, and if $1\leqslant n\leqslant\rho(t)$, then
$\lvert h^{(n)}(t)\rvert<\varepsilon_{*}(t)\cdot\big{(}n\beta(t)\big{)}^{n}\leqslant\varepsilon_{*}(t)\cdot\big{(}\rho(t)\beta(t)\big{)}^{\rho(t)}=\varepsilon(t)$
as required. This finishes the proof of Theorem 1. ∎
Next we prove Corollary 1. Let $I$, $f$, $\varepsilon$, $\rho$ be as in the corollary. The case where $I$ is open is covered by Theorem 1, and the case where $I$ is closed and bounded is a consequence of the Weierstrass Approximation Theorem [19, p. 33]. Suppose $I=[\alpha,\beta)$} where $\infty<\alpha<\beta\leqslant\infty$. (The remaining case where $I=(\alpha,\beta]$ with $\infty\leqslant\alpha<\beta<\infty$ is treated in a similar way.) Then $\operatorname{fr}(I)=\{\beta\}$. We set $I_{1}:=(-\infty,\beta)$ and now use a well-known fact:
Claim : There is a $\mathcal{C}^{r}$-function $f_{1}\colon I_{1}\to{\mathbb{R}}$ which extends $f$.
To see this take $\delta\in{\mathbb{R}}^{>}$ and an extension of $f$ to a $\mathcal{C}^{r}$-function $(\alpha-\delta,\beta)\to{\mathbb{R}}$, also denoted by $f$. Let $\theta\in\mathcal{C}^{\infty}({\mathbb{R}})$ be such that $\theta(t)=0$ for $t\leqslant\alpha-(\delta/2)$ and $\theta(t)=1$ for $t\geqslant\alpha$. (See, e.g., Section 3 below.) Then $f_{1}\colon I_{1}\to{\mathbb{R}}$ given by $f_{1}(t):=f(t)\theta(t)$ if $t\in(\alpha-\delta,\beta)$ and $f_{1}(t):=0$ if $t\leqslant\alpha-\delta$ is $\mathcal{C}^{r}$ and extends $f|_{I}$.
To finish the proof, take $f_{1}$ as in the claim and extend $\varepsilon$, $\rho$ to $\varepsilon_{1},\rho_{1}\in\mathcal{C}(I_{1})$ by $\varepsilon_{1}(t):=\varepsilon(\alpha)$, $\rho_{1}(t):=\rho(\alpha)$ for $t\leqslant\alpha$; then Theorem 1 applied to $f_{1}$$\varepsilon_{1}$$\rho_{1}$ in place of $f$, $\varepsilon$, $\rho$ yields $g_{1}\in\mathcal{C}^{\omega}(I_{1})$ which extends to a holomorphic function ${\mathbb{C}}\setminus\{\beta\}\to{\mathbb{C}}$} such that $\lvert(f_{1}-g_{1})^{(n)}(t)\rvert<\varepsilon_{1}(t)$ for each $t\in I_{1}$ and $n\leqslant\rho_{1}(t)$. Then $g:=g_{1}|_{I}\in\mathcal{C}^{\omega}(I)$ has an extension to a holomorphic function ${\mathbb{C}}\setminus\{\beta\}\to{\mathbb{C}}$} and satisfies $\lvert(f-g)^{(n)}(t)\rvert<\varepsilon(t)$ for $t\in I$ and $n\leqslant\rho(t)$ as required. ∎
3. Bounds on the Derivatives of Bump Functions
By a bump function we mean here a $\mathcal{C}^{\infty}$-function $\alpha\colon{\mathbb{R}}\to{\mathbb{R}}$ which is 0 on $(-\infty,0]$, strictly increasing on $[0,1]$, and $1$ on $1,+\infty)$. Such bump functions play an important role in many constructions in analysis, e.g., partitions of unity. In this section we give a construction of a bump function $\alpha$ with controlled growth of its derivatives $\alpha^{(n)}$. As an auxiliary result we first establish the following formula for the derivatives of reciprocals of nonzero elements of a differential field $K$:
Proposition 3.1.
For $\phi\in K^{\times}$ and $n\geqslant 1$ we have
(3.1) $(\phi^{-1})^{(n)}=\sum_{k=1}^{n}(-1)^{k}{n+1\choose k+1}(\phi^{-1})^{k+1}(\phi^{k})^{(n)}.$
Proof.
We work in the setting of [4, Chapter 12] and assume familiarity with the concepts and the notations introduced there. Let $\phi$ range over $K^{\times}$ and iijjkkll over $\mathbb{N}$}. Put
$h_{\phi}:=\sum_{n}\left(\frac{\phi^{(n)}}{\phi}\right)\frac{z^{n}}{n!}\in 1+zK[[z]]\subseteq K[[z]]^{\times}.$
By the generalized Leibniz Rule, $\phi\mapsto h_{\phi}\colon K^{\times}\to K[[z]]^{\times}$ is a group morphism, so we obtain a group morphism $\phi\mapsto[h_{\phi}]\colon K^{\times}\to\mathfrak{tr}_{K}$ whose image is contained in the Appell group $\mathcal{A}$ over $K$ (cf. [4, p. 565]). For $i\leqslant j$ we have $h_{\phi}]_{ij}={j\choose i}\phi^{(j-i)}/\phi$. Since $\mathcal{A}\subseteq 1+\mathfrak{tr}_{K}^{1}$, we have $\big{(}[h_{\phi}]-1\big{)}^{k}\in\mathfrak{tr}_{K}^{k}$ for $k\geqslant 1$, and
$h_{\phi^{-1}}]=[h_{\phi}]^{-1}=\sum_{k}(-1)^{k}\big{(}[h_{\phi}]-1\big{)}^{k}=\sum_{k}(-1)^{k}\left(\sum_{l=0}^{k}{k\choose l}(-1)^{k-l}[h_{\phi^{l}}]\right).$
For $n\geqslant 1$ this yields
$(\phi^{-1})^{(n)}\phi=[h_{\phi^{-1}}]_{0n}$ $=\sum_{k=1}^{n}(-1)^{k}\left(\big{(}[h_{\phi}]-1\big{)}^{k}\right)_{0n}$
$=\sum_{k=1}^{n}(-1)^{k}\left(\sum_{l=1}^{k}{k\choose l}(-1)^{k-l}[h_{\phi^{l}}]_{0n}\right)$
$=\sum_{l=1}^{n}(-1)^{l}\left(\sum_{k=l}^{n}{k\choose l}\right)(\phi^{-1})^{l}(\phi^{l})^{(n)}$
$=\sum_{l=1}^{n}(-1)^{l}{n+1\choose l+1}(\phi^{-1})^{l}(\phi^{l})^{(n)}$
where for the last equality we used the well-known identity
$\sum_{k=l}^{n}{k\choose l}={n+1\choose l+1}$,
which has an easy proof by induction on $n$. ∎
Remark.
The previous proposition also holds if $K$ is any differential ring (in the sense of [4, 4.1]). To see this let $Y$ be a differential indeterminate over $\mathbb{Q}$}. Then the identity (3.1) holds for ${\mathbb{Q}}\langle Y\rangle$, $Y$ in place of $K$, $\phi$, respectively. It now suffices to note that given any differential ring $K$ and a unit $\phi\in K^{\times}$ we have the morphism $S^{-1}{\mathbb{Q}}\{Y\}\to K$ of differential rings with $Y\mapsto\phi$, where
$S^{-1}{\mathbb{Q}}\{Y\}=\big{\{}Y^{-n}P:P\in{\mathbb{Q}}\{Y\},n\geqslant 0\big{\}}\subseteq{\mathbb{Q}}\langle Y\rangle$
is the localization of ${\mathbb{Q}}\{Y\}$ at its multiplicative subset $S:=\{1,Y,Y^{2},\dots\}$.
In the rest of this section we prove:
Proposition 3.2.
There are a bump function $\alpha$ and constants $c,d\in{\mathbb{R}}^{>}$ such that $\lVert\alpha\rVert_{n}\leqslant c\,n^{dn}$ for each $n\geqslant 1$.
We begin our construction by studying the function $\theta\colon{\mathbb{R}}\to{\mathbb{R}}^{\geqslant}$ given by $\theta(t):=\operatorname{e}^{-1/t}$ if $t>0$ and $\theta(t):=0$ otherwise. For each $n$ and $t>0$ we have
$\theta^{(n)}=\frac{p_{n}(t)}{t^{2n}}\theta(t),$
where $p_{n}\in{\mathbb{R}}[T]$ is the polynomial given recursively by $p_{0}=1$ and
(3.2) $p_{n+1}=T^{2}p_{n}^{\prime}-(2nT-1)p_{n}.$
Thus
$p_{1}=1,\quad p_{2}=-2T+1,\quad p_{3}=6T^{2}-6T+1,\quad p_{4}=-24T^{3}+36T^{2}-12T+1,\quad\dots.$
In general, $\deg p_{n}=n-1$ for $n\geqslant 1$. Since for each $m$ we have $t^{-m}\operatorname{e}^{-1/t}\to 0$ as $t\to 0^{+}$, this yields that $\theta^{(n)}(t)\to 0$ as $t\to 0^{+}$, so $\theta$ is $\mathcal{C}^{\infty}$. We now want to bound the quantities $\lVert\theta^{(n)}\rVert_{[0,1]}$. First, some notation and a lemma.
Notation.
For a polynomial $p=a_{0}+a_{1}T+\cdots+a_{n}T^{n}\in{\mathbb{R}}[T]$ ($a_{0},\dots,a_{n}\in{\mathbb{R}}$) we let $|p|:=\max\big{\{}|a_{0}|,\dots,|a_{n}|\big{\}}$.
Lemma 3.3.
$\lvert p_{n}\rvert\leqslant 2^{n-1}n!$ for $n\geqslant 1$.
Proof.
This is clear for $n=1$. Suppose we have shown the inequality for some $n\geqslant 1$}. Let $a_{0},\dots,a_{n-1},b_{0},\dots,b_{n}\in{\mathbb{R}}$ with $p_{n}=a_{0}+a_{1}T+\cdots+a_{n-1}T^{n-1}$ and $p_{n+1}=b_{0}+b_{1}T+\cdots+b_{n}T^{n}$. Then
$T^{2}p_{n}^{\prime}$ $\ =\ \sum_{m=1}^{n}(m-1)a_{m-1}T^{m},$
$(2nT-1)p_{n}$ $\ =\ -a_{0}+\sum_{m=1}^{n-1}(2na_{m-1}-a_{m})T^{m}+2na_{n}T^{n}.$
Hence by the recursion relation (3.2) we have
$b_{0}=-a_{0},\quad b_{m}=a_{m}+(-2n+m-1)a_{m-1}\quad\text{for $m=1,\dots,n-1$,}\quad b_{n}=-2na_{n-1}$
and this yields $\lvert p_{n+1}\rvert\leqslant 2^{n}(n+1)!$ as required. ∎
Next, given $\lambda\in{\mathbb{R}}^{>}$, consider the function $t\mapsto\zeta(t):=t^{-\lambda}\operatorname{e}^{-1/t}\colon{\mathbb{R}}^{>}\to{\mathbb{R}}^{>}$. Then
$\zeta^{\prime}(t)/\zeta(t)=-\lambda/t+1/t^{2}=(-\lambda+1/t)/t,$
so $\max\zeta({\mathbb{R}}^{>})=\zeta(1/\lambda)=\lambda^{\lambda}\operatorname{e}^{-\lambda}$. Since the function $\lambda\mapsto\lambda^{\lambda}\operatorname{e}^{-\lambda}\colon{\mathbb{R}}^{\geqslant 1}\to{\mathbb{R}}^{>}$ is strictly increasing, this yields, for $t\in[0,1]$ and $n\geqslant 1$:
$\lvert\theta^{(n)}(t)\rvert=\left|\frac{p_{n}(t)}{t^{2n}}\operatorname{e}^{-1/t}\right|\leqslant n\cdot\lvert p_{n}\rvert\cdot(n+1)^{n+1}\operatorname{e}^{-(n+1)}$
and thus
$\log\,\lvert\theta^{(n)}(t)\rvert\leqslant\log n+(n-1)\log 2+n\log n+(n+1)\big{(}\log(n+1)-1\big{)}\leqslant 3n\log n.$
Therefore
(3.3) $\lVert\theta^{(n)}\rVert_{[0,1]}\leqslant n^{3n}\qquad\text{for $n\geqslant 1$.}$
Let $\mu\in{\mathbb{R}}^{>}$. Then $\theta(t)^{\mu}=\theta(t/\mu)$ and thus $(\theta^{\mu})^{(n)}(t)=\theta^{(n)}(t/\mu)/\mu^{n}$, hence
$\lVert(\theta^{\mu})^{(n)}\rVert_{[0,1]}=\mu^{-n}\lVert\theta^{(n)}\rVert_{[0,1/\mu]}\leqslant\lVert\theta^{(n)}\rVert_{[0,1]}\quad\text{if $\mu\geqslant 1$.}$
Also using (3.3), this yields:
(3.4) $\lVert\theta^{\mu}\rVert_{[0,1];\,n}\leqslant n^{3n}\quad\text{if $\mu,n\geqslant 1$.}$
We note that (3.4) also holds with $\theta$ replaced by the $\mathcal{C}^{\infty}$-function $\theta_{*}\colon{\mathbb{R}}\to{\mathbb{R}}^{\geqslant}$ given by $\theta_{*}(t):=\theta(1-t)$. Since $\theta+\theta_{*}>0$, we obtain the $\mathcal{C}^{\infty}$-function
$\alpha:=\theta/(\theta+\theta_{*})\colon{\mathbb{R}}\to{\mathbb{R}}$,
and this is a bump function: we have $\alpha(t)=0$ for $t\leqslant 0$ and $\alpha(t)=1$ for $t\geqslant 1$, and since for $t\in(0,1)$ we also have
$\alpha^{\prime}(t)=\frac{\theta^{\prime}(t)\theta(1-t)+\theta(t)\theta^{\prime}(1-t)}{\big{(}\theta(t)+\theta(1-t)\big{)}^{2}}=\theta(t)\theta(1-t)\frac{t^{-2}+(1-t)^{-2}}{\big{(}\theta(t)+\theta(1-t)\big{)}^{2}}>0,$
the restriction of $\alpha$ to $[0,1]$ is strictly increasing. Our goal is now to compute bounds on $\lVert\alpha\rVert_{n}$, using (3.4). For this we apply the remark after Proposition 3.1 to the differential ring $K=\mathcal{C}^{\infty}({\mathbb{R}})$ and the unit $\phi:=\theta+\theta_{*}$ of $K$. Suppose $n\geqslant 1$. We first note that
$\phi^{k}=\sum_{l=0}^{k}{k\choose l}\theta^{k-l}\theta_{*}^{l}$
where by (0.1) and (3.4) we have
$\lVert\theta^{k-l}\theta_{*}^{l}\rVert_{[0,1];\,n}\leqslant 2^{n}n^{6n}\quad\text{for $l=0,\dots,k$ and $m=0,\dots,n$,}$
hence $\lVert\phi^{k}\rVert_{[0,1];\,n}\leqslant 2^{k+n}n^{6n}$. Next we note that for $t\in(0,1)$ we have
$\phi^{\prime}(t)=\frac{\theta(t)}{t^{2}}-\frac{\theta_{*}(t)}{(1-t)^{2}},$
hence $\phi^{\prime}(t)<0$ if $t<1/2$ and $\phi^{\prime}(t)>0$ if $t>1/2$. This yields $\phi\geqslant\phi(1/2)=\operatorname{e}^{-2}$ and so $\lVert(\phi^{-1})^{k+1}\rVert_{[0,1]}\leqslant\operatorname{e}^{2(k+1)}\leqslant 2^{3(k+1)}$. Using (3.1) we thus conclude:
$\lVert(\phi^{-1})^{(n)}\rVert_{[0,1]}$ $\leqslant\sum_{k=1}^{n}{n+1\choose k+1}\cdot\lVert(\phi^{-1})^{k+1}\rVert_{[0,1]}\cdot\lVert(\phi^{k})^{(n)}\rVert_{[0,1]}$
$\leqslant\sum_{k=1}^{n}{n+1\choose k+1}\cdot 2^{3(k+1)}\cdot 2^{k+n}n^{6n}\leqslant 2^{6n+4}n^{6n}.$
Now $\alpha=\theta\cdot\phi^{-1}$ and thus
$\lVert\alpha\rVert_{n}=\lVert\alpha\rVert_{[0,1];\,n}$ $\leqslant 2^{n}\cdot\lVert\theta\rVert_{[0,1];\,n}\cdot\lVert\phi^{-1}\rVert_{[0,1];\,n}$
$\leqslant 2^{n}\cdot n^{3n}\cdot 2^{6n+4}n^{6n}=2^{7n+4}n^{9n}\leqslant cn^{dn}$
where $c:=2^{11}$, $d:=16$. This concludes the proof of Proposition 3.2. ∎
Remark.
Note that the bound in this proposition is qualitatively optimal in the following sense: there is no function $\gamma\colon{\mathbb{N}}\to{\mathbb{R}}^{>}$ with $\gamma(n)=o(n)$ such that $\lVert\alpha^{(n)}\rVert=O(n^{\gamma(n)})$. To see this suppose $\beta$ is any bump function. Let $n$ be given and put $M:=\lVert\beta^{(n)}\rVert$. If $n\geqslant 1$ then $\lvert\beta^{(n-1)}(t)\rvert\leqslant Mt$ for $t\in(0,1)$, hence if $n\geqslant 2$ then $\lvert\beta^{(n-2)}(t)\rvert\leqslant Mt^{2}/2$ for $t\in(0,1)$, etc., thus $\lvert\beta(t)\rvert\leqslant Mt^{n}/n!$ for $t\in(0,1)$. Hence there is no $c\in{\mathbb{R}}^{>}$ such that $\lVert\beta^{(n)}\rVert\leqslant c^{n}n!$ for all $n$, since for such $c$ we would have $\beta(t)=0$ for each $t\in[0,1/c]$, contradicting that $\beta|_{[0,1]}$ is strictly increasing. Now the claim follows from this observation and Lemma 1.3.
Let $\alpha$, $c$, $d$ be as in Proposition 3.2, where we may assume $c\geqslant 1$. With $0^{0}:=1$, put $C_{n}:=cn^{dn}$, so $1=C_{0}\leqslant C_{1}\leqslant\cdots$. For $a<b$ in $\mathbb{R}$}, we define the increasing $\mathcal{C}^{\infty}$-function $\alpha_{a,b}\colon{\mathbb{R}}\to{\mathbb{R}}$ by
(3.5) $\alpha_{a,b}(t)\ :=\ \alpha\!\left(\frac{t-a}{b-a}\right),$
so $\alpha_{a,b}(t)=0$ for $t\leqslant a$ and $\alpha_{a,b}(t)=1$ for $t\geqslant b$. Also,
$\big{|}\alpha_{a,b}^{(m)}(t)\big{|}\ \leqslant\ \frac{C_{m}}{(b-a)^{m}}\ \text{ for all $m$ and $t$}.$
The case most relevant for us later is when $b-a\leqslant 1$; then we have
(3.6) $\lVert\alpha_{a,b}\rVert_{n}\leqslant\frac{C_{n}}{(b-a)^{n}}.$
Next let $a<b<a_{*}<b_{*}$ be in $\mathbb{R}$}, and with $\varepsilon:=\frac{1}{3}(b-a),\varepsilon_{*}:=\frac{1}{3}(b_{*}-a_{*})\in{\mathbb{R}}^{>}$ define the (hump) function $\alpha_{a,b,a_{*},b_{*}}\colon{\mathbb{R}}\to{\mathbb{R}}$ by
(3.7) $\alpha_{a,b,a_{*},b_{*}}(t)\ :=\ \begin{cases}\alpha_{a+\varepsilon,b-\varepsilon}(t)&\text{if $t\leqslant b$,}\\ 1-\alpha_{a_{*}+\varepsilon_{*},b_{*}-\varepsilon_{*}}(t)&\text{otherwise.}\end{cases}$
Then $\alpha_{a,b,a_{*},b_{*}}(t)=0$ if $t\notin[a,b_{*}]$, $\alpha_{a,b,a_{*},b_{*}}(t)=1$ if $t\in[b,a_{*}]$, and $\alpha_{a,b,a_{*},b_{*}}$ is increasing on $[a,b]$ and decreasing on $a_{*},b_{*}]$. Suppose $\varepsilon,\varepsilon_{*}\leqslant 1$. Then by (3.6),
(3.8) $\lVert\alpha_{a,b,a_{*},b_{*}}\rVert_{n}\ \leqslant\ 3^{n}C_{n}\max\left\{\frac{1}{(b-a)^{n}},\frac{1}{(b_{*}-a_{*})^{n}}\right\}.$
In particular, for suitable reals $c_{1},d_{1}\geqslant 1$ (depending only on $bab_{*}-a_{*}$ccdd) we have $\lVert\alpha_{a,b,a_{*},b_{*}}\rVert_{n}\leqslant c_{1}n^{d_{1}n}$ for each $n$.
4. Weierstrass Transforms
Let $f\in\mathcal{C}^{m}({\mathbb{R}})$ be such that $\operatorname{supp}f$ (= the closure in $\mathbb{R}$} of the set ${\mathbb{R}}\setminus f^{-1}(0)$) is bounded; let also $\lambda$ range over $\mathbb{R}$$^{>}$. The function $f_{\lambda}\colon{\mathbb{R}}\to{\mathbb{R}}$} given by
(4.1) $f_{\lambda}(t):=(\lambda/\pi)^{1/2}\int_{-\infty}^{\infty}f(s)\operatorname{e}^{-\lambda(s-t)^{2}}\,ds$
is known as the (generalized) Weierstrass transform with parameter $\lambda$; below we sometimes denote $f_{\lambda}$ also by $W_{\lambda}(f)$. We could have replaced the bounds -$\infty\infty$ in this integral by any aabb such that $\operatorname{supp}(f)\subseteq[a,b]$. A change of variables gives
$f_{\lambda}(t)\ =\ (\lambda/\pi)^{1/2}\int_{-\infty}^{\infty}f(t-s)\operatorname{e}^{-\lambda s^{2}}\,ds.$
From the Gaussian integral $\int_{-\infty}^{\infty}\operatorname{e}^{-s^{2}}\,ds=\pi^{1/2}$ we get $(\lambda/\pi)^{1/2}\int_{-\infty}^{\infty}\operatorname{e}^{-\lambda s^{2}}\,ds=1$, hence $\lVert f_{\lambda}\rVert\leqslant\lVert f\rVert$. Also, $f_{\lambda}$ extends to an entire function; in particular, $f_{\lambda}\in\mathcal{C}^{\omega}({\mathbb{R}})$. For $k\leqslant m$ we have $(f_{\lambda})^{(k)}=(f^{(k)})_{\lambda}$, so $\lVert f_{\lambda}\rVert_{m}\leqslant\lVert f\rVert_{m}$. See [3, Appendix A] for proofs of these facts. (Entire functions which arise as Weierstrass transforms have been studied extensively [10, 20, 22, 26].) Next we record an explicit version of [3, Lemma A.2], for which we let $f\in\mathcal{C}^{m+1}({\mathbb{R}})$ have bounded support.
Lemma 4.1.
Let $\varepsilon\in{\mathbb{R}}^{>}$, and set
$\lambda(\varepsilon):=8(\lVert f\rVert_{m+1}\varepsilon^{-1})^{2}\log^{+}\!\big{(}2\sqrt{2}\,\lVert f\rVert_{m}\,\varepsilon^{-1}\big{)}\in{\mathbb{R}}^{\geqslant}.$
Then for each $\lambda>\lambda(\varepsilon)$ we have $\lVert f_{\lambda}-f\rVert_{m}\leqslant\varepsilon$. (In particular, $\lVert f_{\lambda}-f\rVert_{m}\leqslant\varepsilon$ provided $\lambda\geqslant 16\sqrt{2}(\lVert f\rVert_{m+1}\varepsilon^{-1})^{3}+1$.)
Proof.
Put $\delta:=(\varepsilon/2)/\lVert f\rVert_{m+1}$. By the Mean Value Theorem we have
$\lvert f^{(k)}(s)-f^{(k)}(t)\rvert\leqslant\varepsilon/2\quad\text{whenever $\lvert s-t\rvert\leqslant\delta$ and $k\leqslant m$.}$
An easy computation shows that $\lambda>\lambda(\varepsilon)$ implies $\sqrt{2}\lVert f\rVert_{m}\operatorname{e}^{-(\lambda/2)\delta^{2}}\leqslant\varepsilon/2$, and by the argument in the proof of Lemma A.2, this guarantees $\lVert f_{\lambda}-f\rVert\leqslant\varepsilon$. ∎
5. Whitney Approximation with Bounds
Let $(a_{n})$, $(b_{n})$, $(\varepsilon_{n})$ be sequences in $\mathbb{R}$} and $(r_{n})$ in $\mathbb{N}$} such that
  1. (i)
    $a_{0}=b_{0}$, $(a_{n})$ is strictly decreasing, $(b_{n})$ is strictly increasing,
  2. (ii)
    $\varepsilon_{n}>\varepsilon_{n+1}$ with $\varepsilon_{n}\to 0$ as $n\to\infty$, and
  3. (iii)
    $r_{n}\leqslant r_{n+1}\leqslant r$ for each $n$.
Set
$K_{n}\ :=\ [a_{n},b_{n}],\qquad L_{n}\ :=\ K_{n+1}\setminus K_{n}\ =\ [a_{n+1},a_{n})\cup(b_{n},b_{n+1}].$
Then $I:=\bigcup_{n}K_{n}$ is an open interval in $\mathbb{R}$}. Let $f\in\mathcal{C}^{r}(I)$. The proof of Whitney’s Approximation Theorem given in [3, Appendix A] then produces a $g\in\mathcal{C}^{\omega}(I)$ such that $\lVert{f-g}\rVert_{L_{n};\,r_{n}}<\varepsilon_{n}$ for each $n$. We recall the main lines of this argument. First, we introduce some hump functions $\varphi_{n}$ as follows: for $n\geqslant 1$, employing the $\mathcal{C}^{\infty}$-functions introduced in (3.7), we set
$\alpha_{n}:=\alpha_{a_{n+2},a_{n+1},a_{n},a_{n-1}},\qquad\beta_{n}:=\alpha_{b_{n-1},b_{n},b_{n+1},b_{n+2}},\qquad\varphi_{n}:=\alpha_{n}+\beta_{n},$
and also set $\varphi_{0}:=\alpha_{a_{2},a_{1},b_{1},b_{2}}\in\mathcal{C}^{\infty}({\mathbb{R}})$. Then for each $n$ we have $\varphi_{n}=0$ on a neighborhood of $K_{n-1}$ (satisfied automatically for $n=0$, by convention), $\varphi_{n}=1$} on a neighborhood of $\operatorname{cl}(L_{n})=[a_{n+1},a_{n}]\cup[b_{n},b_{n+1}]$, and $\operatorname{supp}\varphi_{n}\subseteq K_{n+2}$. With $M_{n}:=1+2^{r_{n}}\lVert\varphi_{n}\rVert_{r_{n}}$, choose $\delta_{n}\in{\mathbb{R}}^{>}$ so that for all $n$,
(5.1) $2\delta_{n+1}\leqslant\delta_{n},\qquad\sum_{m=n}^{\infty}\delta_{m}M_{m+1}\leqslant\varepsilon_{n}/4.$
We inductively define sequences $(\lambda_{n})$ in $\mathbb{R}$$^{>}$ and $(g_{n})$ in $\mathcal{C}^{\omega}({\mathbb{R}})$ as follows: Let $\lambda_{m}\in{\mathbb{R}}^{>}$ and $g_{m}\in\mathcal{C}^{\omega}({\mathbb{R}})$ for $m<n$; then consider $h_{n}\in\mathcal{C}^{r}({\mathbb{R}})$ given by
$h_{n}(t)\ :=\ \begin{cases}\varphi_{n}(t)\cdot\big{(}f(t)-\big{(}g_{0}(t)+\cdots+g_{n-1}(t)\big{)}\big{)}&\text{if $t\in I$,}\\ 0&\text{otherwise.}\end{cases}$
Thus $\operatorname{supp}h_{n}\subseteq\operatorname{supp}\varphi_{n}\subseteq K_{n+2}$ is bounded. Put
$g_{n}:=W_{\lambda_{n}}(h_{n})\in\mathcal{C}^{\omega}({\mathbb{R}})$
where we take $\lambda_{n}\in{\mathbb{R}}^{>}$ such that $\lVert g_{n}-h_{n}\rVert_{r_{n}}<\delta_{n}$: any sufficiently large $\lambda_{n}$ will do, by Lemma 4.1. The argument in [3, Appendix A] then yields a function
$g\colon I\to{\mathbb{R}},\qquad g(t)=\sum_{i=0}^{\infty}g_{i}(t)\text{ for each $t\in I$}$
with $g\in\mathcal{C}^{r_{n}}(I)$ and $\lVert f-g\rVert_{L_{n};\,r_{n}}<\varepsilon_{n}$ for each $n$. Next set
$H_{m}:=2(\lambda_{m}/\pi)^{1/2}\lVert h_{m}\rVert(b_{m+2}-a_{m+2})\in{\mathbb{R}}^{\geqslant},$
and fix a sequence $(c_{m})$ of positive reals such that $M:=\sum_{m\geqslant 1}c_{m}<\infty$. Then we can and do choose the sequences $(g_{m})$, $(\lambda_{m})$ so that in addition
$H_{m}\exp(-\lambda_{m}/m)\ \leqslant\ c_{m}\quad\text{for all $m\geqslant 1$.}$
Therefore
$\sum_{n}H_{n}\exp(-\lambda_{n}\rho)<\infty\qquad\text{for each $\rho\in{\mathbb{R}}^{>}$.}$
Now $g_{n}$ is the restriction to $\mathbb{R}$} of the entire function $\widehat{g}_{n}$ given by
$\widehat{g}_{n}(z)\ =\ (\lambda_{n}/\pi)^{1/2}\int_{a_{n+2}}^{b_{n+2}}h_{n}(s)\operatorname{e}^{-\lambda_{n}(s-z)^{2}}ds.$
In the following we let $x=\operatorname{Re}z$, $y=\operatorname{Im}z$. Then $\operatorname{Re}(s-z)^{2}=(s-x)^{2}-y^{2}$ and so
(5.2) $\lvert\widehat{g}_{n}(z)\rvert\leqslant(\lambda_{n}/\pi)^{1/2}\int_{-\infty}^{\infty}\lvert h_{n}(s)\rvert\operatorname{e}^{-\lambda_{n}\operatorname{Re}(s-z)^{2}}\,ds\leqslant\lVert h_{n}\rVert\operatorname{e}^{\lambda_{n}y^{2}}.$
Put
$\rho_{n}\ :=\ \textstyle\frac{1}{2}\min\!\big{\{}(a_{n}-a_{n+1})^{2},(b_{n+1}-b_{n})^{2}\big{\}}\in{\mathbb{R}}^{>}$
and
$U_{n}\ :=\ \big{\{}z:a_{n+1}<\operatorname{Re}z<b_{n+1},\ \operatorname{Re}\!\big{(}(z-a_{n+1})^{2}\big{)},\ \operatorname{Re}\!\big{(}(z-b_{n+1})^{2}\big{)}>\rho_{n}\big{\}},$
an open subset of $\mathbb{C}$ containing $K_{n}$ such that $\operatorname{Re}\!\big{(}(s-z)^{2}\big{)}>\rho_{n}$ for all $z\in U_{n}$ and $s\in{\mathbb{R}}\setminus K_{n+1}$. If $\rho_{n}\geqslant\rho_{n+1}$, then $U_{n}\subseteq U_{n+1}$. The argument in [3, Appendix A] shows that for each $n$ the series $\sum_{m}\widehat{g}_{m}$ converges uniformly on compact subsets of $U_{n}$, and so we obtain a holomorphic function
$z\mapsto\widehat{g}(z):=\sum_{m}\widehat{g}_{m}(z)\ \colon\ U=\bigcup_{n}U_{n}\to{\mathbb{C}}$
which extends $g$, as claimed. Indeed, if $z\in U_{n}$ and $m\geqslant n+2$, then $\operatorname{supp}h_{m}\subseteq K_{m+2}\setminus K_{m-1}\subseteq{\mathbb{R}}\setminus K_{n+1}$ and so $\lvert\widehat{g}_{m}(z)\rvert\leqslant H_{m}\operatorname{e}^{-\lambda_{m}\rho_{n}}$, hence
(5.3) $\lvert\widehat{g}_{m}(z)\rvert\leqslant H_{m}\operatorname{e}^{-\lambda_{m}/m}\leqslant c_{m}\quad\text{ if $m\geqslant k_{n}:=\max\big{\{}\lceil 1/\rho_{n}\rceil,n+2\big{\}}$}$
and so $\sum_{m\geqslant k_{n}}\lvert\widehat{g}_{m}(z)\rvert\leqslant M$.
In the proofs of Theorems 2 and 3 (given at the end of this section) we need to control the growth of $\lvert\widehat{g}(z)\rvert$ for $z\in U$, and this requires us to choose the quantities $\delta_{n}$$\lambda_{n}$$c_{n}$ in a more explicit way. With this in mind we focus now in particular on the two cases $I={\mathbb{R}}$ and $I={\mathbb{R}}^{>}$, where we assume that $a_{n}$$b_{n}$ are chosen as follows, for some $\delta\in{\mathbb{R}}^{>}$:
  1. ($\mathbb{R}$})
    $a_{n}=-b_{n}=-\delta n$ for each $n$;
  2. ($\mathbb{R}$$^{>}$)
    $a_{n}=\delta/(n+1)$, $b_{n}=\delta(n+1)$ for each $n$.
Below we will label various displayed statements pertaining to these two cases accordingly. For example, note that
(5.4 $\mathbb{R}$}) $\rho_{n}=\delta^{2}/2,\qquad U={\mathbb{C}}.$
This follows by observing that given $z$, for sufficiently large $n$ we have $a_{n+1}=-\delta(n+1)<x<\delta(n+1)=b_{n+1}$ as well as
$(x-a_{n+1})^{2}-y^{2}-\rho_{n}=\big{(}x+\delta(n+1)\big{)}^{2}-y^{2}-(\delta^{2}/2)>0$
and
$(x-b_{n+1})^{2}-y^{2}-\rho_{n}=\big{(}x-\delta(n+1)\big{)}^{2}-y^{2}-(\delta^{2}/2)>0,$
so $z\in U_{n}$. We also have
(5.4 $\mathbb{R}$$^{>}$) $\rho_{n}=\frac{\delta^{2}}{2}\left(\frac{1}{(n+1)(n+2)}\right)^{2},\qquad U=\big{\{}z:\lvert\operatorname{Im}z\rvert<\operatorname{Re}z\big{\}}.$
To see the latter note that if $x>0$ and $y^{2}<x^{2}$, then
$(x-a_{n+1})^{2}-y^{2}-\rho_{n}\to x^{2}-y^{2}>0\text{ as $n\to\infty$,}$
and thus $\operatorname{Re}\big{(}(z-a_{n+1})^{2}\big{)}=(x-a_{n+1})^{2}-y^{2}>\rho_{n}$ for sufficiently large $n$, and likewise, $\operatorname{Re}\big{(}(z-b_{n+1})^{2}\big{)}>\rho_{n}$ for all sufficiently large $n$. Conversely, if $z\in U_{n}$, then $x>a_{n+1}>0$ and $x^{2}-y^{2}\geqslant(x-a_{n+1})^{2}-y^{2}>\rho_{n}>0$.
It will be convenient to assume, in addition to (i)–(iii) above:
  1. (iv)
    $\varepsilon_{n}+\varepsilon_{n+2}\geqslant 2\varepsilon_{n+1}$ for each $n$, and
  2. (v)
    $b_{n}-a_{n}=O(n)$.
(Here (iv) will hold in the settings of the proofs of Theorems 2 and 3, and (v) clearly holds both in the cases ($\mathbb{R}$}) and ($\mathbb{R}$$^{>}$).) We shall also assume $f\neq 0$ below, and let $x=\operatorname{Re}z$, $y=\operatorname{Im}z$.
We now estimate the growth of the derivatives of the $\varphi_{n}$. Let $C_{n}\in{\mathbb{R}}$ be as defined in Section 3. Suppose first we are in the case ($\mathbb{R}$}). By the remark after (3.8) we can take $c,d\in{\mathbb{R}}^{\geqslant 1}$ such that
(5.5 $\mathbb{R}$}) $\lVert\varphi_{n}\rVert_{m}\leqslant c\,m^{dm}\quad\text{for all $m$, $n$.}$
In the case ($\mathbb{R}$$^{>}$) we note $a_{n}-a_{n+1}=\frac{\delta}{(n+1)(n+2)}$ for each $n$, hence for $n\geqslant 1$ we have $\lVert\alpha_{n}\rVert_{m}\leqslant(3/\delta)^{m}C_{m}\big{(}(n+2)(n+3)\big{)}^{m}$ and $\lVert\beta_{n}\rVert_{m}\leqslant(3/\delta)^{m}C_{m}$, so $\lVert\varphi_{n}\rVert_{m}\leqslant(3/\delta)^{m}C_{m}\big{(}(n+2)(n+3)\big{)}^{m}$; we also have $\lVert\varphi_{0}\rVert_{m}\leqslant(18/\delta)^{m}C_{m}$. Thus in this case we obtain $c,d\in{\mathbb{R}}^{>}$ such that
(5.5 $\mathbb{R}$$^{>}$) $\lVert\varphi_{0}\rVert_{m}\leqslant c\,m^{dm},\qquad\lVert\varphi_{n}\rVert_{m}\leqslant c\,(mn)^{dm}\quad\text{if $n\geqslant 1$.}$
Going forward we assume that we have real numbers $D_{mn}$ such that
$D_{0n}=1,\quad 2^{m}\lVert\varphi_{n}\rVert_{m}\leqslant D_{mn}\leqslant D_{m+1,n},D_{m,n+1}\qquad\text{for all $m$, $n$.}$
(In the case ($\mathbb{R}$}), by (5.5 $\mathbb{R}$}), for suitable $c,d\in{\mathbb{R}}^{\geqslant 1}$ we can take $D_{mn}:=c\,m^{dm}$. In the case ($\mathbb{R}$$^{>}$) we may similarly take $D_{m0}:=c\,m^{dm}$ and $D_{mn}:=c\,(mn)^{dm}$ if $n\geqslant 1$, for suitable $c,d\in{\mathbb{R}}^{\geqslant 1}$.) Now put
$D_{n}:=D_{r_{n}n},\qquad N_{n}:=2^{n+1}D_{n}.$
Then $N_{n}\geqslant 2D_{n}$, thus $M_{n}=1+2^{r_{n}}\lVert\varphi_{n}\rVert_{r_{n}}\leqslant 1+D_{n}\leqslant N_{n}$. Also note: $2N_{n}\leqslant N_{n+1}$. We now assume that in the construction of $\widehat{g}$ above we chose
$\delta_{n}:=\frac{\varepsilon_{n}-\varepsilon_{n+1}}{4N_{n+1}}>0.$
Note that these $\delta_{n}$ indeed satisfy the requirements (5.1): we have
$\sum_{m=n}^{\infty}\delta_{m}M_{m+1}\leqslant\sum_{m=n}^{\infty}\delta_{m}N_{m+1}=\frac{1}{4}\sum_{m=n}^{\infty}(\varepsilon_{m}-\varepsilon_{m+1})=\frac{1}{4}\lim_{k\to\infty}(\varepsilon_{n}-\varepsilon_{k+1})\leqslant\frac{1}{4}\varepsilon_{n},$
and we also have $2\delta_{n+1}\leqslant\delta_{n}$ since by (iv):
$\frac{N_{n+2}}{2N_{n+1}}\geqslant 1\geqslant\frac{\varepsilon_{n+1}-\varepsilon_{n+2}}{\varepsilon_{n}-\varepsilon_{n+1}}.$
In order to deduce from Lemma 4.1 a lower bound on just how large $\lambda_{n}$ needs to be taken, we need an upper bound on $\lVert h_{n}\rVert_{m}$ for $m\leqslant r$, which we establish next. For this we note that $\operatorname{supp}\varphi_{n}\subseteq K_{n+2}$ and $\lVert g_{n}\rVert_{m}\leqslant\lVert h_{n}\rVert_{m}$ for each $n$ yields
$\lVert h_{n}\rVert_{m}\leqslant 2^{m}\lVert\varphi_{n}\rVert_{m}\big{(}\lVert f\rVert_{K_{n+2};\,m}+\lVert h_{0}\rVert_{m}+\cdots+\lVert h_{n-1}\rVert_{m}\big{)}.$
An easy induction on $n$ now shows that
$\lVert h_{n}\rVert_{m}\leqslant G_{mn}:=2^{n}(D_{mn})^{n+1}\cdot\lVert f\rVert_{K_{n+2};\,m}.$
In particular $\lVert h_{n}\rVert\leqslant G_{0n}=2^{n}\lVert f\rVert_{K_{n+2}}$. Also note that
$\lVert f\rVert_{K_{n+2}}\leqslant G_{mn}\leqslant G_{m+1,n},G_{m,n+1}.$
Next put
$\mu_{n}:=128\sqrt{2}(\delta_{n}^{-1}G_{r_{n}+1,n})^{3}+1.$
Then $\mu_{n+1}\geqslant\mu_{n}\geqslant 1$, and for each $\lambda\geqslant\mu_{n}$ we have $\lVert{W_{\lambda}(h_{n})-h_{n}}\rVert_{r_{n}}\leqslant\delta_{n}/2<\delta_{n}$ by Lemma 4.1. Below we assume that in our construction of the sequences $(\lambda_{n})$$(g_{n})$ we always chose $\lambda_{n}=\mu_{n}$. Note that $\delta_{n}\leqslant\delta_{0}\leqslant\varepsilon_{0}/2^{n+3}$ and so
$\lambda_{n}\geqslant(\delta_{0}^{-1}G_{0n})^{3}\geqslant 8^{n+3}\varepsilon_{0}^{-3}G_{0n}^{3}.$
Since $f\neq 0$ by assumption, we can take $n_{0}$ with $G_{0n_{0}}>0$. Then with $c:=\varepsilon_{0}^{3}/G_{0n_{0}}^{2}$ we have $c\lambda_{n}\geqslant n^{5}G_{0n}$ for each $n$. The function $t\mapsto t^{2}\operatorname{e}^{-t}\colon{\mathbb{R}}^{>}\to{\mathbb{R}}$ takes on its maximum value $4\operatorname{e}^{-2}<1$ at $t=2$, so $t\operatorname{e}^{-t}<t^{-1}$ for each $t>0$. Hence for $n\geqslant 1$:
$\lambda_{n}^{1/2}\operatorname{e}^{-\lambda_{n}/n}\leqslant n\cdot(\lambda_{n}/n)\operatorname{e}^{-\lambda_{n}/n}\leqslant n^{2}/\lambda_{n}$
and so $\lambda_{n}^{1/2}\operatorname{e}^{-\lambda_{n}/n}G_{0n}\leqslant c/n^{3}$. We have $H_{n}=O\big{(}\lambda_{n}^{1/2}G_{0n}\,(n+2)\big{)}$ by (v), so we obtain a constant $c_{*}>0$ (not depending on the sequence $r_{n}$) such that
$H_{n}\exp(-\lambda_{n}/n)\leqslant c_{*}/n^{2}\quad\text{for $n\geqslant 1$.}$
Hence in the construction of $\widehat{g}$ above we may choose $c_{n}:=c_{*}/n^{2}$ for each $n\geqslant 1$.
Proof of Theorem 2
[…]
$\lambda_{n}=128\sqrt{2}\big{(}\delta_{n}^{-1}G_{r_{n}+1,n}\big{)}^{3}+1$
from above:
Lemma 5.1.
There is a $D\in{\mathbb{R}}^{\geqslant 1}$, independent of $f$, $\varepsilon$, $\rho$, such that for each $n$,
$\lambda_{n}\leqslant\big{(}D(\rho(s)+1)\big{)}^{D(\rho(s)+1)s}\cdot\left(\frac{\lVert f\rVert_{s;\,\rho(s)+1}}{\Delta\varepsilon(s)}\right)^{3}+1\qquad\text{where $s:=\sqrt{2}(n+2)$.}$
Proof.
As we are in case ($\mathbb{R}$}), we can assume that we have $c,d\in{\mathbb{R}}^{\geqslant 1}$ (independent of $f$, $\varepsilon$, $\rho$) such that $D_{mn}=cm^{dm}$ for all mmnn. We have
$\delta_{n}^{-1}=\frac{2^{n+3}D_{n+1}}{\varepsilon_{n}-\varepsilon_{n+1}}\qquad\text{where $D_{n+1}=D_{r_{n+1},n+1}$.}$
Since $\varepsilon$ is convex and decreasing,
$\varepsilon_{n}-\varepsilon_{n+1}=\varepsilon(s-\sqrt{2})-\varepsilon(s)\geqslant\varepsilon(s)-\varepsilon(s+1)=\Delta\varepsilon(s).$
Now $r_{n+1}\leqslant\rho(s)$, thus
$\delta_{n}^{-1}=\frac{c2^{n+3}(r_{n+1})^{dr_{n+1}}}{\varepsilon_{n}-\varepsilon_{n+1}}\leqslant\frac{2c\,2^{s/\sqrt{2}}\rho(s)^{d\rho(s)}}{\Delta\varepsilon(s)}$
and so
$\delta_{n}^{-3}\leqslant\frac{c_{0}\,2^{(3/\sqrt{2})s}\rho(s)^{d_{0}\rho(s)}}{\Delta\varepsilon(s)^{3}}\qquad\text{with $c_{0}:=(2c)^{3}$, $d_{0}:=3d$.}$
Moreover, using that $\rho$ is increasing:
$G_{r_{n}+1,n}$ $=2^{n}(D_{r_{n}+1,n})^{n+1}\cdot\lVert f\rVert_{K_{n+2;\,r_{n}+1}}$
$\leqslant 2^{n}\big{(}c(r_{n}+1)\big{)}^{d(r_{n}+1)(n+1)}\cdot\lVert f\rVert_{K_{n+2;\,r_{n}+1}}$
$\leqslant 2^{s/\sqrt{2}-2}\big{(}c(\rho(s)+1)\big{)}^{d(\rho(s)+1)(s/\sqrt{2}-1)}\cdot\lVert f\rVert_{s;\,\rho(s)+1}$
$\leqslant\big{(}2c(\rho(s)+1)\big{)}^{d(\rho(s)+1)(s/\sqrt{2}-1)}\cdot\lVert f\rVert_{s;\,\rho(s)+1}.$
So with $c_{1}:=2c$, $d_{1}:=3d/\sqrt{2}$ we have
$(G_{r_{n}+1,n})^{3}\leqslant\big{(}c_{1}\big{(}\rho(s)+1\big{)}\big{)}^{d_{1}(\rho(s)+1)s}\cdot\lVert f\rVert_{s;\,\rho(s)+1}^{3}.$
Combining these estimates for $\delta_{n}^{-3}$ and $(G_{r_{n}+1,n})^{3}$ yields $D$ with the required properties. ∎
In the following we set $U_{-1}:=\emptyset$.
Lemma 5.2.
Suppose $z\notin U_{n-1}$. Then $\sqrt{2}n\leqslant\lvert x\rvert+\sqrt{1+y^{2}}\leqslant 1+\sqrt{2}\lvert z\rvert$.
Proof.
The first inequality is clear if $\lvert x\rvert\geqslant\sqrt{2}n$, so suppose $\lvert x\rvert<\sqrt{2}n$. Then $n\geqslant 1$}, and thus $z\notin U_{n-1}$ yields $\operatorname{Re}(z-\sqrt{2}n)^{2}\leqslant 1$ or $\operatorname{Re}(z+\sqrt{2}n)^{2}\leqslant 1$. In the first case $(\sqrt{2}n-x)^{2}\leqslant 1+y^{2}$ where $\sqrt{2}n-x>0$ and thus
$\sqrt{2}n\leqslant x+\sqrt{1+y^{2}}\leqslant\lvert x\rvert+\sqrt{1+y^{2}}.$
In the second case, similarly $(\sqrt{2}n+x)^{2}\leqslant 1+y^{2}$ where $\sqrt{2}n+x>0$ and thus $\sqrt{2}n\leqslant-x+\sqrt{1+y^{2}}\leqslant\lvert x\rvert+\sqrt{1+y^{2}}$. The second inequality is a consequence of Lemma 1.5.∎
Now take $n$ such that $z\in U_{n}\setminus U_{n-1}$. From (5.3) and the remark after it recall that $\sum_{m\geqslant k_{n}}\lvert\widehat{g}_{m}(z)\rvert\leqslant M$ where $k_{n}=n+2$, thus
$\lvert\widehat{g}(z)\rvert\leqslant\sum_{m\leqslant n+1}\lvert\widehat{g}_{m}(z)\rvert+\sum_{m\geqslant n+2}\lvert\widehat{g}_{m}(z)\rvert$
where the second (infinite) sum is $\leqslant M$, so we focus on the first sum. With $s:=\sqrt{2}t+1$ we have $\sqrt{2}n\leqslant s$ by Lemma 5.2, so $n\leqslant s/\sqrt{2}\leqslant t+1\leqslant s$. Recalling that $K_{n+3}=\big{[}{-\sqrt{2}(n+3)},\sqrt{2}(n+3)\big{]}$, we get $G_{0,n+1}=2^{n+1}\lVert f\rVert_{K_{n+3}}\leqslant 2^{s+1}\lVert f\rVert_{s+3\sqrt{2}}$, hence by (5.2), for $m\leqslant n+1$:
$\lvert\widehat{g}_{m}(z)\rvert\leqslant G_{0m}\cdot\exp(\lambda_{m}t^{2})\leqslant G_{0,n+1}\cdot\exp(\lambda_{n+1}t^{2})\leqslant 2^{s+1}\lVert f\rVert_{s+3\sqrt{2}}\cdot\exp(\lambda_{n+1}t^{2}).$
This yields
$\sum_{m\leqslant n+1}\lvert\widehat{g}_{m}(z)\rvert\leqslant(s+2)\cdot 2^{s+1}\lVert f\rVert_{s+3\sqrt{2}}\cdot\exp(\lambda_{n+1}t^{2}),$
hence
$\lVert\widehat{g}\rVert_{t}$ $\leqslant(s+2)\cdot 2^{s+1}\lVert f\rVert_{s+3\sqrt{2}}\cdot\exp(\lambda_{n+1}t^{2})+M$
$\leqslant\exp\big{(}(s+2)+\log^{+}\lVert f\rVert_{s+3\sqrt{2}}+\lambda_{n+1}t^{2}\big{)}+M$
$\leqslant\exp\big{(}N+\log^{+}\lVert f\rVert_{s+3\sqrt{2}}+\lambda_{n+1}s^{2}\big{)}\quad\text{where $N:=\log(1+M)+3$,}$
$\leqslant\exp\big{(}N\cdot s^{2}\cdot(1+\log^{+}\lVert f\rVert_{s+3\sqrt{2}}+\lambda_{n+1})\big{)}.$
Next take $D$ as in Lemma 5.1. Since $\rho$ and $1/\Delta\varepsilon$ are increasing and $s+3\sqrt{2}\geqslant\sqrt{2}(n+3)$, we obtain $\lambda_{n+1}\leqslant\lambda(s+3\sqrt{2})+1$ where
$\lambda(s):=\big{(}D(\rho(s)+1)\big{)}^{D(\rho(s)+1)s}\cdot\left(\frac{\lVert f\rVert_{s;\,\rho(s)+1}}{\Delta\varepsilon(s)}\right)^{3}.$
Hence with $C:=2N$ we have
$\lVert\widehat{g}\rVert_{t}\leqslant\exp\big{(}C\cdot s^{2}\cdot\big{(}1+\log^{+}\lVert f\rVert_{s+3\sqrt{2}}+\lambda(s+3\sqrt{2})\big{)}\big{)}\quad\text{for $t\geqslant 0$, $s=\sqrt{2}t+1$.}$
Here $C$ does not depend on $\rho$, and $D$ does not depend on $f$, $\varepsilon$, $\rho$. This concludes the proof of the theorem. ∎
Proof of Theorem 3
Let $\alpha$, $f$, $\varepsilon$, $\rho$, $V$ be as in Theorem 3, so
$V=\big{\{}z:\operatorname{Re}z>0,\ (\operatorname{Im}z)^{2}\leqslant(\operatorname{Re}z)^{2}-\alpha\big{\}}.$
We choose $(a_{n})$, $(b_{n})$ as in ($\mathbb{R}$$^{>}$) with $\delta:=\sqrt{\alpha/2}$, and we set $\varepsilon_{n}:=\varepsilon(b_{n+1})=\varepsilon\big{(}\delta(n+2)\big{)}$ and $r_{n}:=\lfloor\rho(b_{n+1})\rfloor$ for each $n$. Then (i)–(v) hold, and by (5.4 $\mathbb{R}$$^{>}$) we have $V\subseteq U={\big{\{}z:\lvert\operatorname{Im}z\rvert<\operatorname{Re}z\big{\}}}$ and
$\rho_{n}=\frac{\alpha}{4}\left(\frac{1}{(n+1)(n+2)}\right)^{2}.$
[…]
Lemma 5.3.
Let $z\in V\setminus V_{n-1}$. Then $n\leqslant(2/\sqrt{\alpha})\lvert z\rvert$.
Proof.
We may assume $n\geqslant 1$. Note that $z\in V$ gives $x^{2}\geqslant y^{2}+\alpha\geqslant\alpha$ and $x>0$, so $x\geqslant\sqrt{\alpha}>\delta>a_{n}$. If $x\geqslant b_{n}=\delta(n+1)$, then $n<{n+1}\leqslant(1/\delta)x$, hence $n\leqslant(2/\sqrt{\alpha})\lvert z\rvert$; thus suppose $x<b_{n}$. We first show that $\operatorname{Re}\big{(}{(z-a_{n})^{2}}\big{)}>\rho_{n-1}$. To see this note $z\in V$ also yields $y^{2}\leqslant x^{2}-\alpha$, and together with $a_{n}x<a_{n}b_{n}=\delta^{2}=\alpha/2$,
$\operatorname{Re}\big{(}(z-a_{n})^{2}\big{)}=(x-a_{n})^{2}-y^{2}$ $\geqslant a_{n}^{2}-2a_{n}x+\alpha$
$>a_{n}^{2}=\frac{\delta^{2}}{(n+1)^{2}}$
$\geqslant\frac{\delta^{2}}{2}\left(\frac{1}{n(n+1)}\right)^{2}=\rho_{n-1}$
as claimed. Now $z\notin V_{n-1}$ yields $\operatorname{Re}\big{(}(z-b_{n})^{2}\big{)}\leqslant\rho_{n-1}$, and so
$\big{(}x-\delta(n+1)\big{)}^{2}-y^{2}=\operatorname{Re}\big{(}(z-b_{n})^{2}\big{)}\leqslant\rho_{n-1}\leqslant\alpha/16,$
hence $\delta(n+1)-x\leqslant\sqrt{(\alpha/16)+y^{2}}$. Using Lemma 1.5 we get $\delta(n+1)\leqslant(\sqrt{\alpha}/4)+\sqrt{2}\lvert z\rvert$ and thus $\delta n<\sqrt{2}\lvert z\rvert$, hence $n\leqslant(2/\sqrt{\alpha})\lvert z\rvert$ follows. ∎
Lemma 5.4.
There is a $D\in{\mathbb{R}}^{\geqslant 1}$, independent of $f$, $\varepsilon$, $\rho$, such that for each $n\geqslant 1$,
$\lambda_{n}\leqslant\big{(}D(\rho(s)+1)\big{)}^{D(\rho(s)+1)s^{2}}\cdot\left(\frac{\lVert f\rVert_{s;\,\rho(s)+1}}{\Delta\varepsilon(s)}\right)^{3}+1\qquad\text{where $s:=\delta(n+3)$.}$
Proof.
Since we are in the case ($\mathbb{R}$$^{>}$), we can assume to have $c,d\in{\mathbb{R}}^{\geqslant 1}$ (independent of $f$, $\varepsilon$, $\rho$) such that $D_{mn}=c(mn)^{dm}$ for all $m$ and $n\geqslant 1$. Suppose $n\geqslant 1$, and recall again that
$\delta_{n}^{-1}=\frac{2^{n+3}D_{n+1}}{\varepsilon_{n}-\varepsilon_{n+1}}\qquad\text{where $D_{n+1}=D_{r_{n+1},n+1}$.}$
Using Lemma 1.1, from the convexity of $\varepsilon$ we obtain
$\varepsilon_{n}-\varepsilon_{n+1}=\varepsilon(s-2\delta)-\varepsilon(s-\delta)\geqslant\delta\big{(}\varepsilon(s)-\varepsilon(s+1)\big{)}=\delta\cdot\Delta\varepsilon(s).$
We also have $r_{n+1}\leqslant\rho(b_{n+2})=\rho(s)$ and thus
$\delta_{n}^{-1}\leqslant\frac{c2^{s/\delta}\big{(}\rho(s)(s/\delta-2)\big{)}^{d\rho(s)}}{\varepsilon_{n}-\varepsilon_{n+1}}\leqslant\frac{(c/\delta)\,2^{s/\delta}\,\big{(}\rho(s)(s/\delta)\big{)}^{d\rho(s)}}{\Delta\varepsilon(s)},$
so
$\delta_{n}^{-3}\leqslant\frac{c_{0}2^{(3/\delta)s}\big{(}\rho(s)(s/\delta)\big{)}^{d_{0}\rho(s)}}{\Delta\varepsilon(s)^{3}}\qquad\text{with $c_{0}:=(c/\delta)^{3}$, $d_{0}:=3d$.}$
Moreover,
$G_{r_{n}+1,n}$ $=2^{n}(D_{r_{n}+1,n})^{n+1}\cdot\lVert f\rVert_{K_{n+2};\,r_{n}+1}$
$\leqslant 2^{n}\big{(}c(r_{n}+1)n\big{)}^{d(r_{n}+1)(n+1)}\cdot\lVert f\rVert_{K_{n+2};\,r_{n}+1}$
$\leqslant\big{(}2c(\rho(s)+1)\big{)}^{d(\rho(s)+1)(s/\delta-3)(s/\delta-2)}\cdot\lVert f\rVert_{s;\,\rho(s)+1}$
and thus
$(G_{r_{n}+1,n})^{3}\leqslant\big{(}c_{1}\big{(}\rho(s)+1\big{)}\big{)}^{d_{1}(\rho(s)+1)s^{2}}\cdot\lVert f\rVert_{s;\,\rho(s)+1}^{3}$
where $c_{1}:=2c$, $d_{1}:=3d/\delta^{2}$, and this yields the lemma. ∎
Recall that
$k_{n}=\max\big{\{}\lceil 1/\rho_{n}\rceil,n+2\big{\}}\leqslant\max\left\{(4/\alpha)\big{(}(n+1)(n+2)\big{)}^{2},n+1\right\}+1,$
so we obtain a constant $c_{0}\in{\mathbb{R}}^{>}$, only depending on $\alpha$, such that
$k_{n}+2\leqslant c_{0}(t^{2}+1)\qquad\text{for all $n$ and $t\geqslant(\delta/\sqrt{2})n$.}$
Let now $z\in V$ with $\lvert z\rvert\leqslant t$; we aim to estimate $\lvert\widehat{g}(z)\rvert$ in terms of $t$. Take $n$ with $z\in V_{n}\setminus V_{n-1}$, so $t\geqslant(\delta/\sqrt{2})n$ by Lemma 5.3. We have $\sum_{m\geqslant k_{n}}\lvert\widehat{g}(z)\rvert\leqslant M$ by (5.3). Also $K_{m+2}=\big{[}\delta/(m+3),\delta(m+3)\big{]}$ and so $G_{0m}=2^{m}\lVert f\rVert_{K_{m+2}}\leqslant 2^{m}\lVert f\rVert_{\delta(m+3)}$, hence if $m<k_{n}$ then $G_{0m}\leqslant 2^{c_{0}(t^{2}+1)}\lVert f\rVert_{c_{1}(t^{2}+1)}$ where $c_{1}:=\delta c_{0}$. This yields
$\sum_{m<k_{n}}\lvert\widehat{g}_{m}(z)\rvert$ $\leqslant k_{n}\cdot 2^{c_{0}(t^{2}+1)}\lVert f\rVert_{\delta c_{0}(t^{2}+1)}\cdot\exp(\lambda_{k_{n}-1}t^{2})$
$\leqslant c_{0}(t^{2}+1)\cdot 2^{c_{0}(t^{2}+1)}\lVert f\rVert_{c_{1}(t^{2}+1)}\cdot\exp(\lambda_{k_{n}-1}t^{2}),$
and hence for suitable $N\in{\mathbb{R}}^{>}$, not depending on $\rho$:
$\lvert\widehat{g}(z)\rvert$ $\leqslant\sum_{m<k_{n}}\lVert\widehat{g}_{m}(z)\rVert+\sum_{m\geqslant k_{n}}\lVert\widehat{g}_{m}(z)\rVert$
$\leqslant c_{0}(t^{2}+1)\cdot 2^{c_{0}(t^{2}+1)}\lVert f\rVert_{\delta c_{0}(t^{2}+1)}\cdot\exp(\lambda_{k_{n}-1}t^{2})+M$
$\leqslant\exp\big{(}N\cdot(t^{2}+1)\cdot(1+\log^{+}\lVert f\rVert_{c_{1}(t^{2}+1)}+\lambda_{k_{n}-1})\big{)}.$
Now put $C:=2N/c_{1}$, take $D$ as in Lemma 5.4, and let
$s:=c_{1}(t^{2}+1),\qquad\lambda(s):=\big{(}D(\rho(s)+1)\big{)}^{D(\rho(s)+1)s^{2}}\cdot\left(\frac{\lVert f\rVert_{s;\,\rho(s)+1}}{\Delta\varepsilon(s)}\right)^{3}.$
Then $\lambda_{k_{n}-1}\leqslant\lambda(s)+1$, and so $\lvert\widehat{g}(z)\rvert\leqslant\exp\big{(}C\cdot s\cdot\big{(}1+\log^{+}\lVert f\rVert_{s}+\lambda(s)\big{)}\big{)}$. Here $C$ does not depend on $\rho$, and $c_{1}$, $D$ only on $\alpha$. This concludes the proof of Theorem 3. ∎